### v1

# EFX Exists for Three Types of Agents

Abstract
In this paper, we study the problem of finding an envy-free allocation of indivisible goods among multiple agents. EFX, which stands for envy-freeness up to any good, is a well-studied relaxation of the envy-free allocation problem and has been shown to exist for specific scenarios. For instance, EFX is known to exist when there are only three agents [efx_3], and for any number of agents when there are only two types of valuations [mahara].
In this paper, we show that EFX allocations exist for any number of agents when there are at most three types of additive valuations.
1 Introduction
Fair division of indivisible resources is a well-researched problem at the intersection of theoretical computer science and economics. The problem arises in a variety of practical settings, from allocating slots or assets to distributing aid or shared goods. One of the most intuitive notions of fairness is envy-freeness (EF) [foley1967resource], where each individual is content with their share compared to others. However, when the resources are indivisible - such as physical objects, housing units, or assets like artwork — achieving true envy-freeness becomes impossible in many cases. While EF provides a natural measure of fairness, the combinatorial nature of indivisible goods often renders EF allocations unattainable, highlighting the necessity for more nuanced fairness criteria.
One such refinement is Envy-Freeness up to any Good (EFX), proposed by Caragiannis et al. [efx_no1], which has gained prominence due to its minimal relaxation of EF. Under EFX, an agent may envy another, but such envy can always be resolved by removing any single good from the envied agent’s bundle. Despite its apparent simplicity, the existence of EFX allocations for an arbitrary number of agents remains an unresolved question. This is true even in many restricted settings, where only partial results are known for specific cases [efx_no1, efx_3, mahara, pr20, ghosal2023efxexistsagentstypes, BergerCFF22].
To discuss EFX allocations in more detail, let us consider the formal framework for discrete fair division. Here, we have a set $\mathcal{N}=\{a_{1},a_{2},\ldots,a_{n}\}$ of $n$ agents and a set =$\mathcal{M}=\{g_{1},g_{2},\ldots,g_{m}\}$ of $m$ indivisible goods. Each agent $a_{i}\in\mathcal{N}$ has a valuation function $v_{i}:2^{\mathcal{N}}\to\mathbb{R}_{\geq 0}$ that quantifies the amount of utility derived by agent $a_{i}$ from any given subset of goods. We focus on additive valuations, which imply that the total value derived by a subset of goods is simply the sum of the values of the individual goods within that subset. Formally, for any subset $S\subseteq\mathcal{M}$, the valuation function can be expressed as: $v_{i}(S)=\sum_{g\in S}v_{i}(\{g\})$. An allocation $X=\langle X_{a_{1}},X_{a_{2}},\ldots,X_{a_{n}}\rangle$ is a partition of the goods into $n$ parts, called as bundles, such that agent $a_{i}$ receives bundle $X_{a_{i}}$.
EFX Allocations:
Given an allocation $X$, an agent $a_{i}$ envies agent $a_{j}$ if $v_{i}(X_{a_{i}})<v_{i}(X_{a_{j}})$. This envy is called weak envy if it disappears after removing any single good from $X_{a_{j}}$. Otherwise, the envy is considered strong. An allocation is said to be envy-free up to any good (EFX) if no agent has strong envy towards any other agent. It is not known whether EFX allocations always exist for $n$ agents with arbitrary additive valuations. However, over the years, several partial results have emerged. Following its introduction by Caragiannis et al. [efx_no1], Plaut and Roughgarden [pr20] showed that EFX allocations exist in two specific cases: $(i)$ when there are only two agents, and $(ii)$ for any number of agents, provided all agents have identical valuations. Mahara [mahara] extended this to show that EFX allocation always exists when each agent is of one of two given types, where agents of the same type have identical additive valuations. In a breakthrough paper, Chaudhury et al. [efx_3] showed that EFX allocations always exist for instances with three agents. Ghosal et al. [ghosal2023efxexistsagentstypes] extended this result to $n$ agents when $n-2$ of them have identical valuations.
1.1 Our Contributions
The existence of EFX allocations for three types of agents has been an open problem (see, e.g., [Mahara21]). In this work, we resolve this by providing a constructive proof of the following theorem.
Theorem 1.
EFX allocation always exists for $n$ agents when there are three types of agents, where agents of the same type have identical additive valuations.
1.2 Related Work
The literature on the fair division of indivisible goods is extensive. This section highlights the works most relevant to our study - EFX allocations of goods; for further details, we refer the reader to a recent survey by Amanatidis et al. [amanatidis2023fair].
In addition to the studies on complete EFX allocations discussed above, there have been several results on almost complete EFX allocations, where the objective is to minimize the number of unallocated goods. A notable contribution by Chaudhury et al. [chaudhury2021little] show that EFX allocations exist with at most $n-1$ unallocated goods, ensuring that no agent envies the set of unallocated goods. Following this, Berger et al. [BergerCFF22] and Mahara [Mahara21] show that EFX allocations can be achieved with at most $n-2$ unallocated goods. Moreover, they show that for four agents, EFX allocations can always be computed with at most one unallocated good.
An interesting constraint is to restrict the number of possible values agents can have for the goods. Amanatidis et al. [amanatidis2021maximum] show that EFX allocations exist for bi-valued instances, where each agent assigns one of two possible values to each good. Garg and Murhekar [garg2023computing] extended this result by showing Pareto optimal EFX allocations always exists for bi-valued instances. In a recent article, Amanatidis et al. [amanatidis2024pushing] show that EFX allocations exist for tri-valued instances.
There have been several more studies on approximate notions of EFX allocations. An allocation is said to be $\alpha$-EFX, if for every pair of agents $a_{i}$ and $a_{j}$, agent $a_{i}$ finds her bundle $X_{a_{i}}$ to be at least as valuable as $\alpha\cdot v_{a_{i}}(X_{a_{j}}\setminus g)$ for any $g\in X_{a_{j}}$. Chan et al. [chan2019maximin] showed the existence of $\frac{1}{2}$-EFX allocations. Following this, Amanatidis et al. [AMANATIDIS202094] showed that $0.618$-EFX allocations always exists. In a recent work, Amanatidis et al. [amanatidis2024pushing] show that $\frac{2}{3}$-EFX allocations exist when there are at most seven agents.
2 Preliminaries
Let $\mathcal{N}=\{a_{1},a_{2},\dots,a_{n}\}$ be a set of $n$ agents and =$\mathcal{M}=\{g_{1},g_{2},\dots,g_{m}\}$ be a set of $m$ indivisible goods. A discrete fair division instance is represented by the tuple $\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$, where $\mathcal{V}=\langle v_{a_{1}},v_{a_{2}},\dots,v_{a_{n}}\rangle$ comprises the valuation functions. For each agent $a_{i}\in\mathcal{N}$, the function $v_{i}:2^{\mathcal{M}}\to\mathbb{R}_{{\geq}0}$ specifies the utility that agent $a_{i}$ derives from any given subset of goods.
We use the term bundle to denote a subset of goods. An allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$ is a tuple of $n$ mutually disjoint bundles such that bundle $X_{a_{i}}$ is assigned to agent $a_{i}$ for all $a_{i}\in\mathcal{N}$. We say an allocation $X$ is complete if $\bigcup_{a_{i}\in\mathcal{N}}X_{a_{i}}=\mathcal{M}$.
Let $a_{i}\in\mathcal{N};\ g\in\mathcal{M};\ S,T\subseteq\mathcal{M}$ and $v_{a_{i}}$ be the valuation function of $a_{i}$. To simplify notation, we write $v_{a_{i}}(g)$ to denote $v_{a_{i}}(\{g\})$ and use $S\setminus g$, $S\cup g$ to denote $S\setminus\{g\}$, $S\cup\{g\}$, respectively. We also write $S>_{a_{i}}T$ to denote $v_{a_{i}}(S)>v_{a_{i}}(T)$ and similarly for $<_{a_{i}},\geq_{a_{i}},\leq_{a_{i}}$ and $=_{a_{i}}$. We use $\min_{a_{i}}(S,T)$ and $\max_{a_{i}}(S,T)$ to denote $\operatorname*{arg\,min}_{Y\in\{S,T\}}v_{a_{i}}(Y)$ and $\operatorname*{arg\,max}_{Y\in\{S,T\}}v_{a_{i}}(Y)$ respectively.
Definition 1 (Pareto domination).
An allocation $Y$ Pareto dominates an allocation $X$, denoted by $Y\succ X$, If for every agent $a_{i}\in\mathcal{N}$ her bundle in $Y$ is at least as good as her bundle in $X$ and for at least one agent, her bundle in $Y$ is strictly better than the one in $X$. Formally, $Y\succ X$ if
$\forall a_{i}\in\mathcal{N},Y_{a_{i}}\geq_{a_{i}}X_{a_{i}}\quad\text{and}\quad\exists a_{j}\in\mathcal{N},Y_{a_{j}}>_{a_{j}}X_{a_{j}}$
Given an allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$, we say that agent $a_{i}$ envies another agent $a_{j}$ if $X_{a_{j}}>_{a_{i}}X_{a_{i}}$. We often also say that agent $a_{i}$ envies the bundle $X_{a_{j}}$. For an allocation $X$, the envy relation among the agents is graphically represented as follows.
Definition 2 (Envy Graph).
Given an allocation $X$, let $E_{X}=(\mathcal{N},E)$ be a directed graph where the set of agents $\mathcal{N}$ is the set of vertices and for every pair $(a_{i},a_{j})$ of agents, $(a_{i},a_{j})\in E$ iff agent $a_{i}$ envies agent $a_{j}$ under the allocation $X$.
The envy from an agent can be further categorized into strong and weak envy as follows.
Definition 3 (Strong envy).
Given an allocation $X$, an agent $a_{i}$ strongly envies an agent $a_{j}$ if $X_{a_{i}}<_{a_{i}}X_{a_{j}}\setminus g$ for some $g\in X_{a_{j}}$.
An envy that is not strong is called as a weak envy. Given the definitions of strong and weak envies, an EFX allocation can be defined as follows.
Definition 4 (EFX).
An allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$ is said to be envy free upto any good (EFX) if no agent strongly envies any other agent. That is:
$\forall a_{i},a_{j}\in\mathcal{N},\forall g\in X_{a_{j}},\quad X_{a_{i}}\geq_{a_{i}}X_{a_{j}}\setminus g.$
Non-degenerate instances [efx_3, efx_3_simple]
An instance $\mathcal{I}=\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$ is said to be non-degenerate if and only if no agent values two different bundles equally. That is, $\forall{a_{i}\in\mathcal{N}}$ we have $v_{a_{i}}(S)\neq v_{a_{i}}(T)$ for all $S\neq T,\text{ where }S,T\subseteq\mathcal{M}$. Akrami et al. [efx_3_simple] showed that it suffices to deal with non-degenerate instances when there are $n$ agents with general valuation functions, i.e., if each non-degenerate instance has an EFX allocation, each general instance has an EFX allocation.
In the rest of the paper, we only consider non-degenerate instances. This implies that all goods are positively valued by all agents as value of the empty bundle is assumed to be zero. Non-degenerate instances have the following property.
Proposition 1.
Given any allocation $X$, for any two agents $a$ and $a^{\prime}$ with identical valuation, either $a$ envies $a^{\prime}$ or vice versa.
If an agent $a_{i}$ envies a bundle $S$, we can define a minimally envied subset as follows.
Definition 5 (Minimally envied subset [ChaudhuryKMS21]).
Given an allocation $X$, if an agent $a_{i}$ with bundle $X_{a_{i}}$ envies a bundle $S$, we call $T\subseteq S$ a minimally envied subset of $S$ for agent $a_{i}$ if both the following conditions hold.
  1. 1.
    $X_{a_{i}}<_{a_{i}}T$
  2. 2.
    $X_{a_{i}}\geq_{a_{i}}T\setminus{h}\quad\forall{h\in T}$
For an agent $a_{i}$ and two subsets $S,T\subseteq\mathcal{M}$ such that $S<_{a_{i}}T$, we define $R$ as the minimally envied subset of $T$ with respect to $S$, denoted $R=\mathsf{MES}_{a_{i}}(T,S)$, if $R$ is the subset of $T$ that would be minimally envied by $a_{i}$ assuming $S$ is agent $a_{i}$’s bundle. When $S$ is same as $X_{a_{i}}$, we simply refer to $\mathsf{MES}_{a_{i}}(T,S)$ as $\mathsf{MES}_{a_{i}}(T)$.
Definition 6 (Champion [chaudhury2021little]).
Let $X$ be an allocation and $T\subseteq\mathcal{M}$ be a bundle such that at least one agent envies $T$. An agent $a_{i}$ is said to champion the bundle $T$ if no agent strongly envies the minimally envied subset $\mathsf{MES}_{a_{i}}(T,X_{a_{i}})$.
Definition 7 ($g$-champion [chaudhury2021little]).
Given a partial allocation $X$ and an unallocated good $g$, an agent $a_{i}$ is said to $g$-champion agent $a_{j}$ if $a_{i}$ champions the bundle $X_{a_{j}}\cup g$. If $a_{i}=a_{j}$, we say agent $a_{i}$ self $g$-champions.
Similar to the envy graph $E_{X}$, we can define a champion graph with respect to an unallocated good as follows.
Definition 8 (Champion Graph).
Given a partial allocation $X$ along with an unallocated good $g$, a champion graph, denoted by $M_{X}=(\mathcal{N},E_{M})$, is a directed graph with the set of agents $\mathcal{N}$ as the vertices and for every pair of agents $a_{i},a_{j}$, the edge $(a_{i},a_{j})\in E_{M}$ if $a_{i}$ $g$-champions $a_{j}$.
Throughout this paper, “champion edges” refers to $g$-champion edges for some unallocated good $g$, unless otherwise specified. The following is a simple observation about the champion graphs.
Proposition 2.
For any partial allocation $X$ and corresponding champion graph $M_{X}$ with respect to an unallocated good $g$, $M_{X}$ is always cyclic.
Proof.
Recall that every agent values every good non-zero. For each agent $a_{i}$, as $a_{i}$ envies $X_{a_{i}}\cup g$ over $X_{a_{i}}$, some agent (possibly $a_{i}$) must $g$-champion $a_{i}$. Thus, every agent has an incoming $g$-champion edge, making $M_{X}$ cyclic. ∎
Pareto Improvements:
Given an EFX allocation $X$, which is not necessarily complete, several techniques can be applied to find new EFX allocations that Pareto dominate $X$. We will discuss some of these techniques below.
Firstly, we can eliminate envy cycles.
Lemma 1 ([lipton]).
If $X$ is an EFX allocation such that the envy graph $E_{X}$ of $X$ has a directed cycle, then there exists another EFX allocation $Y$ such that $Y\succ X$.
Chaudhury et al. generalized this idea as follows.
Lemma 2 ([chaudhury2021little]).
Let $X$ be a partial EFX allocation. If there exists two agents $a_{i}$ and $a_{j}$ such that $a_{i}$ $g$-champions $a_{j}$ and $a_{i}$ is reachable from $a_{j}$ in the envy graph $E_{X}$, then there exists another EFX allocation $Y$ such that $Y\succ X$. Moreover, all the agents in the path from $a_{i}$ to $a_{j}$, including $a_{i}$ and $a_{j}$, get strictly better bundles in $Y$.
As a corollary of the above lemma, we can make the following two observations.
Corollary 1 ([chaudhury2021little]).
Let $X$ be a partial EFX allocation. If at least one of the following holds, then there exists another EFX allocation $Y$ such that $Y\succ X$.
  1. 1.
    Envy graph $E_{X}$ has a single source vertex.
  2. 2.
    There exists an agent $a_{i}$ who self $g$-champions.
Given an allocation $X$, if agent $a_{i}$ does not envy agent $a_{j}$ but $a_{i}$ $g$-champions $a_{j}$, it follows that $g\in\mathsf{MES}_{a_{i}}(X_{a_{j}}\cup g,X_{a_{i}})$. In this case, we say that agent $a_{i}$ decomposes the bundle $X_{a_{j}}$ into top and bottom halves. The top half is defined as $T_{a_{j}}^{a_{i}}=\mathsf{MES}_{a_{i}}(X_{a_{j}}\cup g,X_{a_{i}})\setminus g$, while the bottom half is given by $B_{a_{j}}^{a_{i}}=X_{a_{j}}\setminus T_{a_{j}}^{a_{i}}$. This decomposition allows us to compare the top halves of different agents’ bundles, leading to the following observation.
Proposition 3.
Consider an allocation $X$. Suppose agent $a_{i}$ $g$-champions agent $a_{j}$ and does not $g$-champion $a_{\ell}$. Suppose agent $a_{k}$, for $k\neq i$, $g$-champions agent $a_{\ell}$. The bundle $X_{a_{i}}$ is decomposed as $T_{a_{j}}^{a_{i}}$ and $B_{a_{j}}^{a_{i}}$. Similarly, $X_{a_{\ell}}$ is decomposed into $T_{a_{\ell}}^{a_{k}}$ and $B_{a_{\ell}}^{a_{k}}$. Then $T_{a_{j}}^{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}$.
Proof.
Agent $a_{i}$ $g$-champions $a_{j}$ but not $a_{\ell}$. Therefore, $T_{a_{j}}^{a_{i}}\cup g>_{a_{i}}X_{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}\cup g$, which implies $T_{a_{j}}^{a_{i}}>_{a_{i}}T_{a_{\ell}}^{a_{k}}$. ∎
2.1 Three Types of Agents
Given an instance $\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$ of discrete fair division, we say that the agents are of three types if the set of agents can be partitioned as $\mathcal{N}=\mathbb{A}\uplus\mathbb{B}\uplus\mathbb{C}$, where agents in each of the sets $\mathbb{A}$, $\mathbb{B}$, and $\mathbb{C}$ have identical valuation functions. We refer to these as the three groups of agents. In particular, we denote the agents in group $\mathbb{A}$ as $\mathbb{A}=\{a_{1},a_{2},a_{3},\dots,a_{p}\}$, all of whom have the same additive valuation function $v_{a}$. Similarly, the agents in group $\mathbb{B}=\{b_{1},b_{2},\dots,b_{q}\}$ and group =$\mathbb{C}=\{c_{1},c_{2},\dots,c_{r}\}$ have identical additive valuation functions $v_{b}$ and $v_{c}$, respectively. Let $p,q$ and $r$ be the number of agents in groups $\mathbb{A},\mathbb{B}$ and $\mathbb{C}$ respectively. Thus, $|\mathcal{N}|=n=p+q+r$.
Given an allocation $X$ for a non-degenerate instance with three types of agents, for any two agents $a_{i}$ and $a_{j}$ of the same type, we know that either $X_{a_{i}}>X_{a_{j}}$ or $X_{a_{i}}<X_{a_{j}}$. Therefore, w.l.o.g. we assume that the bundles are sorted among the agents in each group and therefore the agents $a_{1},b_{1}$ and $c_{1}$ have the least valued bundles of their respective groups. We call them the leading agents, denoted by $L=\{a_{1},b_{1},c_{1}\}$. Throughout this paper, we will maintain this property as an invariant for each EFX allocation.
Definition 9 (Ordering Invariant ).
Let $X$ be an EFX allocation. For any group $T\in\{\mathbb{A},\mathbb{B},\mathbb{C}\}$ with valuation $v_{t}$, if $\exists i>j$ such that $v_{t}(X_{t}^{i})<v_{t}(X_{t}^{j})$ then we interchange the bundles of the two agents. Thus, we maintain the invariant that $v_{t}(X_{t}^{i})\geq v_{t}(X_{t}^{j})$ whenever $i>j$, for each group $T$ with valuation function $v_{t}$. In particular, the leading agent within each group has the minimum valued bundle within that group.
For an instance with three types of agents, Proposition 2 can be refined as follows.
Proposition 4.
Given an instance with $3$ types of agents, and an allocation $X$, only leading agents can be the sources in the envy graph $E_{X}$. Hence $E_{X}$ has at most $3$ sources. Moreover, the following hold:
  1. 1.
    If any non-leading agent $x$ envies (respectively $g$-champions) another agent $y$, then the leading agent in the group of $x$ also envies (respectively $g$-champions) $y$.
  2. 2.
    In the champion graph $M_{X}$, there must be a cycle of length at most $3$, and the cycle involves only the leading agents.
Proof.
By the Ordering Invariant (Definition 9), if $x$ envies $y$, the leading agent in the same group as $x$ has the same valuation function and a smaller valued bundle than $x$. Hence Part 1 follows.
To see Part 2, consider allocation of $g$ to each of the leading agents. From Proposition 4, we know that $M_{X}$ is cyclic. If a leading agent self-$g$-champions, then $M_{X}$ has a cycle of length $1$. If no leading agent self-$g$-champions, then by Part 1, no agent in any group $g$-champions the leading agent in the same group. So the leading agent of each group must be $g$-championed by a leading agent of another group, thereby creating a $2$-cycle or a $3$-cycle. ∎
2.2 Progress Measures
A common approach to proving the existence of EFX allocations is to begin with an empty, trivially EFX allocation, and iteratively extend it. To ensure termination, various progress measures have been proposed. Pareto dominance is one such well-studied measure, where each partial EFX allocation Pareto dominates the previous one. Chaudhury et al. [efx_3] introduced the lexicographic improvement to prove the existence of EFX for three agents.
To prove Theorem 1, we use a combination of Pareto dominance and a new potential function as defined below.
Definition 10 (Potential function: ).
For any, possibly partial, allocation $X$, potential of $X$ is defined to be
$\phi(X)=\min_{a_{i}\in\mathbb{A}}v_{a}(X_{a_{i}})$
Note that the function $\phi$ is compatible with Pareto dominance. That is,
Proposition 5.
For allocations $X$, $Y$, if $Y\succ X$, then $\phi(Y)\geq\phi(X)$.
If $\phi(Y)>\phi(X)$, we say $Y$ $\phi$-dominates $X$. We start with a trivially EFX allocation and iteratively find new EFX allocations that either Pareto dominate or $\phi$-dominate the previous one. As Pareto dominance is a partial order on the set of allocations, and $\phi$ forms a total order, this procedure must terminate.
Given a partial EFX allocation $X$, we discuss a few simple methods to obtain a $\phi$-dominating EFX allocation.
Firstly, if agent $a_{1}$ is the only source in the envy graph $E_{X}$, then as a corollary of Lemma 2, we get a $\phi$-dominating EFX allocation. This is true as the bundle value of agent $a_{1}$ strictly increases.
Corollary 2.
Given a partial EFX allocation $X$, if $a_{1}$ is the only source in $E_{X}$, then there exists an EFX allocation $Y$, possibly partial, such that $\phi(Y)>\phi(X)$.
We generalize this idea to cases where there is no actual cycle among the agents, but rather among the groups. We call this the $\phi$-improving exchanges, as discussed in the following section.
2.3 $\phi$-improving exchanges
Similar to Pareto Improving Cycles in [BergerCFF22], we define $\phi$-improving exchanges for instances with three types of agents. These are not necessarily cycles among the agents, but rather cycles among the groups. Some examples of a $\phi$-improving exchanges are given in Figure 1.
Definition 11 ($\phi$-improving exchange).
[…]
(a) $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ is a $\phi$-improving exchange involving two compatible champion edges.
[…]
(b) $\langle(a_{1},b_{i}),(b_{1},c_{1}),(c_{1},a_{1})\rangle$ is a $\phi$-improving exchange involving one $g$-champion edge
[…]
(c) $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ is a $\phi$-improving exchange with only envy edges
[…]
(d) $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ is a $\phi$-improving exchange
Figure 1: Some examples of $\phi$-improving exchanges
Given an EFX allocation $X$, possibly partial, a $\phi$-improving exchange involving three groups is a tuple of edges either of the form $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ or $\langle(a_{1},c_{j}),(c_{1},b_{i}),(b_{1},a_{1})\rangle$, where $b_{i}\in\mathbb{B}$ and $c_{j}\in\mathbb{C}$, such that the following holds:
  1. (i)
    every edge in the tuple is either an envy edge or a $g$-champion edge for some unallocated good $g$.
  2. (ii)
    For any unallocated good $g$, the tuple can have at most one $g$-champion edge.
Similarly, a $\phi$-improving exchange involving two groups is a tuple of the form $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ (or $\langle(a_{1},c_{j}),(c_{1},a_{1})\rangle$) such that the above conditions hold.
To perform a $\phi$-improving exchange, within each group, the bundle that is being envied (or championed) is given to the envying (or championing) group. Within each group, the incoming bundle is then allocated to the agent who was being envied (or championed) in the exchange. If the edge is a champion edge, the $\mathsf{MES}$ bundle is transferred instead of the whole bundle.
Performing a $\phi$-exchange on an EFX allocation $X$ immediately gives a $\phi$-dominant EFX allocation $Y$. This is stated in the following Lemma 3.
Lemma 3.
A $\phi$-improving exchange on an EFX allocation $X$ gives another EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
Proof.
Consider an EFX allocation $X$. A $\phi$-improving exchange in $X$ can either involve agents from two groups, or from all the three groups.
If it involves two groups, say $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$, we form a new allocation $Y$ by assigning $X_{a_{1}}$ to $b_{i}$ (or $\mathsf{MES}_{b}(X_{a_{1}}\cup g)$ if $(a_{1},b_{i})$ is a $g$-champion edge) and $X_{b_{i}}$ (or $\mathsf{MES}_{a}(X_{b_{i}}\cup h)$ for an $h$-champion edge) to $a_{i}$.
If it involves three groups, w.l.o.g let $\langle(a_{1},b_{i}),(b_{1},c_{j}),(c_{1},a_{1})\rangle$ be a $\phi$-improving exchange. The new allocation $Y$ is obtained by transferring the bundle of $b_{i}$ to $a_{1}$, that of $c_{j}$ to $b_{1}$, and the bundle of $a_{1}$ to $c_{1}$. If any of these are champion edges, we allocate $\mathsf{MES}$ bundles accordingly.
The bundles are then re-sorted to restore the Ordering Invariant. Let $X(A)$ be the set of bundles of agents in $\mathbb{A}$ in the allocation $X$. In allocation $Y$, the value of the minimum valued bundle in $\mathbb{A}$ increases since the new bundles in $\mathbb{A}$ i.e. $Y(A)=(X(A)\setminus\{X_{a_{1}})\}\cup\{X_{b_{i}}\}$, and $v_{a}(X_{b_{i}})>v_{a}(X_{a_{1}})$. So the new minimum valued bundle in $\mathbb{A}$ is $\min\{X_{a_{2}},X_{b_{i}}\}$.
Now we prove that $Y$ is an EFX allocation. For each agent, the bundle in $Y$ is either the same as that in $X$, or a better bundle than that in $X$. Also, the exchanged bundles are either the same bundles as in $X$ or are an $\mathsf{MES}$ of $g$ added to a bundle in $X$. Therefore, the exchanged bundles are not strongly envied by any agent. This proves that $Y$ is also an EFX allocation. ∎
2.4 Proof outline
To prove Theorem 1, we show that, for any partial EFX allocation $X$, there is another EFX allocation $Y$ such that either $Y\succ X$ or $\phi(Y)>\phi(X)$. Since the number of allocations is finite, this process must terminate with a complete EFX allocation. The proof of Theorem 1 is organized into the following case configurations, where in each configuration, we show how to construct the allocation $Y$.
  1. 1.
    Agent $a_{1}$ is a source in $E_{X}$: This is further subdivided into the following sub-cases:
    1. (a)
      Agent $a_{1}$ envies someone in $\mathbb{B}\cup\mathbb{C}$. This configuration is considered in Lemma 7 and Lemma 8 in Section 4.
    2. (b)
      Agent $a_{1}$ envies no one in $\mathbb{B}\cup\mathbb{C}$. The existence of $Y$ is proved in Lemma 9 in Section 5.
  2. 2.
    Agent $a_{1}$ is not a source in $E_{X}$: This case is considered in Section 6. Two sub-cases arise, as mentioned below:
    1. (a)
      Agent $a_{1}$ envies some agent in $\mathbb{B}\cup\mathbb{C}$. In this case, the construction of the allocation $Y$ is shown in Lemma 14.
    2. (b)
      Agent $a_{1}$ does not envy any agent in $\mathbb{B}\cup\mathbb{C}$. This case is considered in Lemma 15.
Key recurring configurations.
In proving these cases, several key configurations arise frequently, which help in constructing the desired allocation $Y$.
First, we show that when there are three sources in the envy graph $E_{X}$, we can obtain $Y$,$Y\succ X$. This is shown in Lemma 6 of Section 3.
In many scenarios, we encounter allocations such that the strong envies are confined within groups, meaning there is no inter-group strong envy. We show that such envy can be resolved without reducing the potential $\phi$. This is shown more generally in the following lemma.
Lemma 4 (Resolving internal strong envies).
Let $X$ be any allocation. Suppose an agent, say $a_{i}$, strongly envies a set $S$ of agents only within her group, i.e., $S\subset\mathbb{A}$. Then, there exists another allocation $Y$ such that the following conditions hold:
  1. (i)
    $\phi(Y)=\phi(X)$
  2. (ii)
    In $Y$, agent $a_{i}$ does not strongly envy anyone.
  3. (iii)
    In $Y$, no agent in $S$ strongly envies any agent.
Proof.
Let $a_{j}\in S$. So $a_{i}$ strongly envies $a_{j}$. Then we replace $X_{a_{j}}$ by $Y_{a_{j}}=\mathsf{MES}_{a_{i}}(X_{a_{j}})$. Clearly $X_{a_{j}}$ is not the minimum valued bundle in $\mathbb{A}$. So $\phi$ doesn’t change by this replacement. Since such a replacement is done for each $a_{j}\in S$, $a_{i}$ has no strong envy towards any agent in $S$. Since $a_{i}$ did not strongly envy anyone in other groups with respect to $X$, and since all the agents outside $S$ have the same bundles in $X$ and $Y$, $a_{i}$ does not strongly envy anyone outside $S$ either. Although agents in $S$, like $a_{j}$, got smaller bundles, $Y_{a_{j}}>_{a}X_{a_{i}}$. So $a_{j}$ does not strongly envy anyone with respect to $Y$ as $a_{i}$ does not strongly envy anyone with respect to $Y$.∎
Since Lemma 4 applies to any of the three groups, we have the following corollary.
Corollary 3.
If $X$ is an allocation where no agent strongly envies another agent from a different group, then there exists an EFX allocation $Y$,possibly partial, such that $\phi(Y)=\phi(X)$.
There are cases where two of the leading agents champion each other, while a third agent envies neither of them, where the following lemma applies.
Lemma 5.
Let $X$ be a partial EFX allocation with an unallocated good $g$ such that in $X$, two of the leading agents (say, $b_{1}$ and $c_{1}$) $g$-champion each other and none of the leading agents envy those two agents ($b_{1}$ and $c_{1}$, in this case). Then, there exists an EFX allocation $Y$ such that $Y\succ X$. Moreover, the bundles of both the leading agents in the $g$-champion cycle are strictly improved.
Proof.
Without loss of generality, assume that agents $b_{1}$ and $c_{1}$ $g$-champion each other. Then, the bundles of $b_{1}$ and $c_{1}$ can be partitioned into top and bottom parts as follows:
[…]
We consider a new allocation by exchanging the two top halves as follows:
[…]
From Proposition 3, we know that $T_{b_{1}}^{c_{1}}>_{c_{1}}T_{c_{1}}^{b_{1}}$ and $T_{c_{1}}^{b_{1}}>_{b_{1}}T_{b_{1}}^{c_{1}}$. Therefore, both agents $b_{1}$ and $c_{1}$ are strictly better off. Now, consider the potential strong envies in $X^{\prime}$. Since the allocation $X$ was EFX, any strong envies in $X^{\prime}$ must be towards $b_{1}$ or $c_{1}$. As $T_{b_{1}}^{c_{1}}>_{c_{1}}T_{c_{1}}^{b_{1}}$, we know that
$X^{\prime}_{c_{1}}=T_{b_{1}}^{c_{1}}\cup B_{c_{1}}^{b_{1}}>T_{c_{1}}^{b_{1}}\cup B_{c_{1}}^{b_{1}}=X_{c_{1}}>T_{b_{1}}^{c_{1}}\cup B_{b_{1}}^{c_{1}}=X_{b_{1}}>T_{c_{1}}^{b_{1}}\cup B_{b_{1}}^{c_{1}}=X^{\prime}_{b_{1}}$
Therefore, no agent from group $\mathbb{C}$ envies agent $b_{1}$. Similarly, no one from $\mathbb{B}$ envies agent $c_{1}$. So, the only possible strong envies are from agents in $\mathbb{A}$ to either $2$ or $3$. We know that if any agent in $\mathbb{A}$ envies some agent $a^{*}\in\mathcal{N}$, then $a_{1}$ also envies $a^{*}$. Since agent $a_{1}$ envied neither $b_{1}$ nor $c_{1}$ in the allocation $X$, agent $a_{1}$ can envy at most one agent among $b_{1}$ and $c_{1}$ in $X^{\prime}$. Without loss of generality, assume that agent $a_{1}$ strongly envies agent $b_{1}$. Then, we replace $B_{b_{1}}^{c_{1}}$ in agent $b_{1}$’s bundle with $g$ to get the following allocation:
[…]
As $T_{c_{1}}^{b_{1}}\cup g$ is a minimally envied subset, no agent strongly envies $b_{1}$. Since $T_{c_{1}}^{b_{1}}\cup g>_{b_{1}}X_{b_{1}}$, she does not envy agent $c_{1}$. Furthermore, since agent $c_{1}$ did not self $g$-champion in $X$, no agent in $\mathbb{C}$ envy agent $b_{1}$. Finally, if any agent from $\mathbb{C}\setminus{c_{1}}$ strongly envies $c_{1}$, we replace $c_{1}$’s bundle with $\mathsf{MES}_{\mathbb{C}}(Y_{c_{1}},X_{c_{2}})$, which is still better than $X_{c_{1}}$ for agent $c_{1}$. Therefore, $Y$ is an EFX allocation such that $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus{L})=X(\mathcal{N}\setminus{L})$.∎
3 Three sources in $E_{X}$
As the leading agents hold the minimum valued bundle in their respective groups, they envy all the agents from their own group and are never envied by anyone in their own group. We consider the case when no leading agent is envied by anyone. That is, $a_{1},b_{1}$ and $c_{1}$ are the sources in $E_{X}$.
Lemma 6.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If the set of leading agents $L=\{a_{1},b_{1},c_{1}\}$ are sources in the envy graph $E_{X}$, then there exists an allocation $Y$ such that $Y\succ X$. Furthermore, $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus{L})=X(\mathcal{N}\setminus{L})$.
Proof overview:
In this case, we only update the bundles of the leading agents, while the remaining agents retain their old bundles. After the update, each leading agent prefers their new bundle over their old ones. Additionally, the leading agents do not envy the new bundles of other leading agents, even when compared to their own old bundles. Thus, the non-leading agents in a group (who keep their old bundles) do not strongly envy any agent in other groups. As a result, the only strong envies that remain are within the groups, which are resolved using Lemma 4. Therefore, the new allocation we obtain is an EFX allocation. Furthermore, the new bundles satisfy $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus L)=X(\mathcal{N}\setminus L)$.
Proof.
From Proposition 4, we know that the $g$-champion graph $M_{X}$ must contain a cycle of length at most three, involving agents $a_{1},b_{1}$, and $c_{1}$. If any of the leading agents self $g$-champions, using Corollary 1, we get a Pareto-dominating EFX allocation $Y$. If the $g$-champion cycle is of length $2$, then by applying Lemma 5, we get our desired allocation. Therefore, assume that $M_{X}$ has a cycle of length $3$. Without loss of generality, assume that $(a_{1},c_{1})$,$(c_{1},b_{1})$ and $(b_{1},a_{1})$ are $g$-champion edges in $M_{X}$ as shown in Figure 2. Since these agents do not envy each other, the $g$-champion edges break the bundle into top and bottom halves as shown below.
[…]
For ease of notation, we will call $T_{a_{1}}^{b_{1}},T_{b_{1}}^{c_{1}}$, and $T_{c_{1}}^{a_{1}}$ as $T_{a},T_{b}$ and $T_{c}$. Similarly, we call the bottom halves as $B_{a},B_{b}$ and $B_{c}$. We now make the following observations regarding the relationships among these parts of the bundles.
Firstly, agent $a_{1}$ champions $c_{1}$, but does not self $g$-champion. Therefore, $T_{c}\cup g>_{a}T_{a}\cup g\implies T_{c}>_{a}T_{a}$. Similarly, as $a_{1}$ does not $g$-champion $b_{1}$, we have $T_{c}\cup g>_{a}T_{b}\cup g\implies T_{c}>_{a}T_{b}$.
As $T_{c}>_{a}T_{a}$, we have $T_{c}\cup B_{a}>_{a}T_{a}\cup B_{a}=X_{a_{1}}$. As $a_{1}$ does not envy $c_{1}$, we know that $X_{a_{1}}=T_{a}\cup B_{a}>_{a}T_{c}\cup B_{c}$. By combining these two inequalities we get $T_{c}\cup B_{a}>_{a}T_{c}\cup B_{c}\implies B_{a}>_{a}B_{c}$.
As $a_{1}$ does not self $g$-champion, we have $T_{a}\cup B_{a}>_{a}T_{a}\cup g\implies B_{a}>_{a}g$. As $a_{1}$ $g$-champions $c_{1}$ but does not envy $c_{1}$, we have $T_{c}\cup g>_{a}T_{c}\cup B_{c}\implies g>_{a}B_{c}$.
Analogous observations can be made for agents $b_{1}$ and $c_{1}$. The following table summarizes the above observations for all three leading agents:
Agent Observations
$a_{1}:\quad$ $T_{c}>_{a}T_{a};\quad T_{c}>_{a}T_{b};\quad B_{a}>_{a}g>_{a}B_{c}$ (1)
$b_{1}:\quad$ $T_{a}>_{b}T_{b};\quad T_{a}>_{b}T_{c};\quad B_{b}>_{b}g>_{b}B_{a}$ (2)
$c_{1}:\quad$ $T_{b}>_{c}T_{c};\quad T_{b}>_{c}T_{a};\quad B_{c}>_{c}g>_{c}B_{b}$ (3)
Furthermore, as $a_{1}$ does not envy $b_{1}$, we have $T_{a}\cup B_{a}>_{a}T_{b}\cup B_{b}$. Therefore, if $B_{a}<_{a}B_{b}$, then it must be the case that $T_{a}>_{a}T_{b}$. Combining with Condition 1, if $B_{a}<_{a}B_{b}$, then $T_{c}>_{a}T_{a}>_{a}T_{b}$ and $B_{b}>_{a}B_{a}>_{a}g>_{a}B_{c}$. Analogous observations can be made for agents $b_{1}$ and $c_{1}$.
We have three predicates: $B_{a}<_{a}B_{b},B_{b}<_{b}B_{c}$, and $B_{c}<_{c}B_{a}$. We have the following four possibilities based of how many of these predicates hold true.
  1. Case 1
    All the predicates hold: That is, $B_{a}<_{a}B_{b},B_{b}<_{b}B_{c}$, and $B_{c}<_{c}B_{a}$. We then have the following updated inequalities:
    Agent Observations
    $a_{1}:\quad$ $T_{c}>_{a}T_{a}>_{a}T_{b}\quad\text{and}\quad B_{b}>_{a}B_{a}>_{a}g>_{a}B_{c}$ (4)
    $b_{1}:\quad$ $T_{a}>_{b}T_{b}>_{b}T_{c}\quad\text{and}\quad B_{c}>_{b}B_{b}>_{b}g>_{b}B_{a}$ (5)
    $c_{1}:\quad$ $T_{b}>_{c}T_{c}>_{c}T_{a}\quad\text{and}\quad B_{a}>_{c}B_{c}>_{c}g>_{c}B_{b}$ (6)
    We allocate $T_{c}\cup B_{b}$ to agent $a_{1}$, $T_{a}\cup B_{c}$ to agent $b_{1}$, and $T_{b}\cup B_{a}$ to agent $c_{1}$. These bundles correspond to the most preferred top and bottom halves for each agent (refer to Conditions 4, 5, and 6). Therefore, all three agents strictly improve. Furthermore, according to agent $a_{1}$, we have $T_{c}\cup B_{b}>_{a}T_{a}\cup B_{a}=X_{a_{1}}>_{a}T_{a}\cup B_{c}=Y_{b_{1}}$. That is, $a_{1}$ does not envy $b_{1}$ even with respect to her old bundle $X_{a_{1}}$. Therefore, no agent in $\mathbb{A}$ envy $b_{1}$. Similarly, we have $X_{a_{1}}>_{a}T_{b}\cup B_{a}=Y_{b_{1}}$. That is, no agent from $\mathbb{A}$ envies $b_{1}$. Similarly we can show that no agents from either $\mathbb{B}$ or $C$ envy any of $a_{1},b_{1}$ or $c_{1}$. If any strong envy remains, it must be confined to agents within their respective groups. Hence, we apply Lemma 4 to resolve these envies.
  2. Case 2
    None of the three predicates holds: Thus $B_{a}>_{a}B_{b}$, $B_{b}>_{b}B_{c}$, and $B_{c}>_{c}B_{a}$. We then cyclically shift the top halves of these three bundles, allocating $T_{c}\cup B_{a}$ to agent $a_{1}$, $T_{a}\cup B_{b}$ to agent $b_{1}$, and $T_{b}\cup B_{c}$ to agent $c_{1}$. Since $T_{c}>_{a}T_{a}$ and $T_{c}>_{a}T_{b}$, agent $a_{1}$ strictly improves and does not envy either $b_{1}$ or $c_{1}$ as they received less preferred top halves according to $a_{1}$. Similarly, both $b_{1}$ and $c_{1}$ also experience improvements and do not envy any agents from $\{a_{1},b_{1},c_{1}\}$. Furthermore, by applying the inequalities 1,2, and 3 we have $T_{b}>_{a}T_{a}$ which implies $T_{b}\cup B_{b}>_{a}T_{a}\cup B_{b}=Y_{b_{1}}$. Along with the fact that $a_{1}$ did not previously envy $b_{1}$, we have $X_{a_{1}}>_{a}X_{b_{1}}>_{a}Y_{b_{1}}$. Therefore agent $a_{1}$ does not envy $b_{1}$ even with respect to her old bundle $X_{a_{1}}$. Therefore, no agent in $\mathbb{A}$ envy $b_{1}$. In an analogous way we can show that none of the agents envy any of $a_{1},b_{1}$ or $c_{1}$. If any strong envy remains, it must be confined to agents within their respective groups. Hence, we apply Lemma 4 to resolve these envies.
  3. Case 3
    One of the three predicates holds: Without loss of generality, assume $B_{a}<_{a}B_{b}$, $B_{b}>_{b}B_{c}$, and $B_{c}>_{c}B_{a}$.
    We then have:
    Agent Observations
    $a_{1}:\quad$ $T_{c}>_{a}T_{a}>_{a}T_{b}\quad\text{and}\quad B_{b}>_{a}B_{a}>_{a}g>_{a}B_{c}$ (4)
    $b_{1}:\quad$ $T_{a}>_{b}T_{b};\quad T_{a}>_{b}T_{c};\quad B_{b}>_{b}g>_{b}B_{a}$ (2)
    $c_{1}:\quad$ $T_{b}>_{c}T_{c};\quad T_{b}>_{c}T_{a};\quad B_{c}>_{c}g>_{c}B_{b}$ (3)
    We allocate the bundles as: $T_{c}\cup B_{a}$ to agent $a_{1}$, $T_{a}\cup g$ to $b_{1}$ and $T_{b}\cup B_{c}$ to agent $c_{1}$.
    Firstly, for agent $a_{1}$, we have $T_{c}\cup B_{a}>_{a}T_{a}\cup B_{a}$ and $T_{a}\cup B_{a}>_{a}T_{a}\cup g$. Therefore, agent $a_{1}$ strictly improves and does not envy $b_{1}$ even with respect to her old bundle. Thus no agent in $\mathbb{A}$ envy $b_{1}$. Furthermore, since $T_{c}\cup B_{a}>_{a}T_{a}\cup B_{a}>_{a}T_{b}\cup B_{c}$, neither agent $a_{1}$ nor any other agent from $\mathbb{A}$ envies $c_{1}$.
    Secondly, as $T_{a}\cup g$ is the $\mathsf{MES}$ bundle, agent $b_{1}$ strictly improves, i.e., $T_{a}\cup g>_{b}T_{b}\cup B_{b}$. Given the case assumption $B_{b}>_{b}B_{c}$, we have $T_{b}\cup B_{b}>_{b}T_{b}\cup B_{c}$. Thus, none of the agents from $\mathbb{B}$ envy $c_{1}$. As $b_{1}$ did not $g$-champion $c_{1}$, we have $T_{b}\cup B_{b}>_{b}T_{c}\cup g$. From Condition 2, we know that $T_{c}\cup g>_{b}T_{c}\cup B_{a}$. Therefore, as $T_{b}\cup B_{b}>_{b}T_{c}\cup B_{a}$, no agent in $\mathbb{B}$ envies $a_{1}$.
    Finally, As $T_{b}>_{c}T_{c}$, agent $c_{1}$ also has strict improvement. That is, $T_{b}\cup B_{c}>_{c}T_{c}\cup B_{c}$. Because of the case assumption $B_{c}>_{c}B_{a}$, we have $T_{c}\cup B_{c}>_{c}T_{c}\cup B_{a}$. As no agent in $\mathbb{C}$ previously envied $X_{c_{1}}=T_{c}\cup B_{c}$, no agent in $\mathbb{C}$ envies $a_{1}$ in $Y$. As $c_{1}$ or any other agent in $\mathbb{C}$ did not previously $g$-champion $a_{1}$, they do not envy $b_{1}$ who has the bundle $T_{a}\cup g$.
    Therefore, if any strong envy remains, it must be confined to agents within their respective groups. Hence, we apply Lemma 4 to resolve these envies.
  4. Case 4
    Two out of the three predicates hold true: Without loss of generality, assume $B_{a}<_{a}B_{b}$, $B_{b}<_{b}B_{c}$, and $B_{c}>_{c}B_{a}$. We have the following observations:
    Agent Observations
    $a_{1}:\quad$ $T_{c}>_{a}T_{a}>_{a}T_{b}\quad\text{and}\quad B_{b}>_{a}B_{a}>_{a}g>_{a}B_{c}$ (4)
    $b_{1}:\quad$ $T_{a}>_{b}T_{b}>_{b}T_{c}\quad\text{and}\quad B_{c}>_{b}B_{b}>_{b}g>_{b}B_{a}$ (5)
    1. Case 4.1
      $\mathbf{T_{c}>_{c}T_{a}}$: Now, along with the above two observations, we have
      $c_{1}:\quad$ $T_{b}>_{c}T_{c}>_{c}T_{a}\quad\text{and}\quad B_{c}>_{c}g>_{c}B_{b}\quad\text{and}\quad B_{c}>_{c}B_{a}$ (6)
      We allocate $T_{c}\cup B_{b}$ to $a_{1}$, $T_{a}\cup B_{c}$ to $b_{1}$ and $T_{b}\cup g$ to agent $c_{1}$.
      For agents $a_{1}$ and $b_{1}$, it is clear from Conditions 4 and 5 that they have received strict improvements with respect to their previous bundles and that they do not envy anyone in $\{a_{1},b_{1},c_{1}\}$ in the new allocation. Furthermore, from 4 we see that $T_{a}\cup B_{a}>_{a}T_{a}\cup B_{c}=Y_{b_{1}}$ and $T_{a}\cup B_{a}>_{a}T_{b}\cup g=Y_{c_{1}}$. Therefore, no agent from $\mathbb{A}$ envies either $b_{1}$ or $c_{1}$. Similarly, using 5 we can show that agents in $\mathbb{B}$ envy neither $a_{1}$ nor $c_{1}$.
      For agent $c_{1}$, we know that $T_{b}\cup g>_{c}T_{c}\cup B_{c}$ as $c_{1}$ $g$-championed $b_{1}$. By applying Condition 6, we get $T_{c}\cup B_{c}>_{c}T_{c}\cup B_{b}$ and $T_{c}\cup B_{c}>_{c}T_{a}\cup B_{c}$. Therefore, agent $c_{1}$ (or any other agent in $\mathbb{C}$) envies neither $b_{1}$ nor $a_{1}$. Therefore, any remaining strong envy must be restricted to agents within their respective groups. Consequently, we apply Lemma 4 to resolve these envies.
    2. Case 4.2
      $\mathbf{T_{c}<_{c}T_{a}}$: Therefore, instead of Condition 6, we have the following observation:
      $c_{1}:\quad$ $T_{b}>_{c}T_{a}>_{c}T_{c}\text{ and }B_{c}>_{c}g>_{c}B_{b}\text{ and }B_{c}>_{c}B_{a}$ (7)
      In this case, we would like to allocate $T_{c}\cup B_{a}$ to $a_{1}$, $T_{b}\cup B_{b}$ to $b_{1}$ and $T_{a}\cup B_{c}$ to agent $c_{1}$. However, in this allocation agent $b_{1}$ envies agent $c_{1}$, and this envy could potentially be a strong envy. Therefore, we consider the $\mathsf{MES}$ bundle $S=\mathsf{MES}_{b}(T_{a}\cup B_{c},X_{b_{1}})$. Based on whether or not agent $c_{1}$ prefers $S$ over $X_{c_{1}}$, we have the following two sub-cases:
      1. Case 4.2(a)
        $\mathbf{S>_{c}X_{c_{1}}}$: Then, we allocate the bundles as follows. Agent $a_{1}$ gets $T_{c}\cup B_{a}$, $b_{1}$ retains $T_{b}\cup B_{b}$ and $c_{1}$ gets $S$. From Condition 4 we have $T_{c}\cup B_{a}>_{a}T_{a}\cup\mathbb{B}_{a}$ and from this case assumtion we know that $S>_{c}X_{c_{1}}$. Therefore agents $a_{1}$ and $c_{1}$ receive strict improvements while agent $b_{1}$ retains her old bundle.
        We know that $X_{a_{1}}>_{a}X_{b_{1}}=Y_{b_{1}}$ Therefore, not agent from $\mathbb{A}$ envies $b_{1}$. Furthermore, as $B_{a}>_{a}B_{c}$ (Condition 4), we get $T_{a}\cup B_{a}>_{a}T_{a}\cup B_{c}>_{a}S$. Thus, no agent from $\mathbb{A}$ envies $c_{1}$.
        The bundle $S$ is an $\mathsf{MES}$ bundle of agent $b_{1}$. Therefore, agent $b_{1}$ (or any other agent in $\mathbb{B}$) does not strongly envy $c_{1}$. From Condition 5, we have $T_{a}>_{b}T_{c}$. Therefore, $T_{a}\cup B_{a}>_{b}T_{c}\cup B_{a}$. Since $b_{1}$ did not previously envy $a_{1}$, we also have $T_{b}\cup B_{b}>_{b}T_{a}\cup B_{a}$. By combining these two inequalities, we get $T_{b}\cup B_{b}>_{b}T_{c}\cup B_{a}$. Hence, $b_{1}$ (and therefore no one else from $\mathbb{B}$) does not envy $a_{1}$.
        We know that $S>_{c}X_{c_{1}}=T_{c}\cup B_{c}$. From Case $4$ assumption we have $B_{c}>_{c}B_{a}$. Therefore, $S>_{c}T_{c}\cup B_{c}>_{a}T_{c}\cup B_{a}$. Therefore, none of the agents in $\mathbb{C}$ envy $a_{1}$. As $c_{1}$ or any other agent in $\mathbb{C}$ did not previously envy $b_{1}$, they do not envy $b_{1}$ now.
        Therefore, any remaining strong envy must be restricted to agents within their respective groups. Consequently, we apply Lemma 4 to resolve these envies.
      2. Case 4.2(b)
        $\mathbf{S<_{c}X_{c_{1}}}$: Then, we allocate the bundles as follows: Agent $a_{1}$ gets $T_{c}\cup B_{b}$, $b_{1}$ gets $S$ and $c_{1}$ gets her $\mathsf{MES}$ bundle $T_{b}\cup g$.
        According to agents in $\mathbb{A}$: Applying Condition 4, $T_{c}\cup B_{b}>_{a}T_{a}\cup B_{a}>_{a}T_{a}\cup B_{c}>_{a}S$. Therefore, $a_{1}$’s bundle strictly improved and $a_{1}$ does not envy $b_{1}$ even with respect to her old bundle. Thus no agent in $\mathbb{A}$ envy $b_{1}$. Since $a_{1}$ or anyone else from $\mathbb{A}$ did not previously $g$-champion $b_{1}$, They do not envy $c_{1}$ who has the bundle $T_{b}\cup g$.
        According to agents in $\mathbb{B}$: As $S$ is $b_{1}$’s $\mathsf{MES}$ bundle, $b_{1}$ received a strictly better bundle. From Condition 5, we know that $T_{b}>_{b}T_{c}$, therefore, $S>_{b}T_{b}\cup B_{b}>_{b}T_{c}\cup B_{b}$. That is, $b_{1}$ does not envy $a_{1}$, even with respect to her old bundle. Thus, no agent in $\mathbb{B}$ envy $a_{1}$. Since $b_{1}$ did not previously self $g$-champion, and none of the agents in $\mathbb{B}$ $g$-championed $b_{1}$, they do not envy $T_{b}\cup g$ of agent $c_{1}$.
        According to agents in $\mathbb{C}$: $T_{b}\cup g$ is an $\mathsf{MES}$ bundle of agent $c_{1}$. Therefore, $c_{1}$ received strict improvement. Using Condition 7, we know $B_{c}>_{c}B_{b}$. Therefore, $T_{c}\cup B_{c}>_{c}T_{c}\cup B_{b}$. That is, $c_{1}$ does not envy $a_{1}$, even with respect to her old bundle. Thus no agent in $\mathbb{C}$ envies $a_{1}$ From the current case assumption, that is $S<_{c}X_{c_{1}}$, we know that no agent in $\mathbb{C}$ envies $b_{1}$.
        Therefore, the current allocation Pareto dominates $X$ and any strong envy that remains must be restricted to agents within their respective groups. Thus, we apply Lemma 4 to resolve these envies. ∎
Corollary 4.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If the set of leading agents $L=\{a_{1},b_{1},c_{1}\}$ are all the sources in the envy graph $E_{X}$, then there exists an EFX allocation $Y$, possibly partial, such that at least one of the following holds:
  1. 1.
    $\phi(Y)>\phi(X)$
  2. 2.
    $Y\succ X$ and $a_{1}$ is a source in $E_{Y}$ and $\forall\alpha\in(\mathbb{B}\cup\mathbb{C})\setminus L$ if $(a_{1},\alpha)$ is an edge in $E_{X}$ then $(a_{1},\alpha)$ is an edge in $E_{Y}$.
Proof.
Given a partial EFX allocation $X$ with three sources, we apply Lemma 6 to get an allocation $Y$ such that $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus{L})=X(\mathcal{N}\setminus{L})$. If agent $a_{1}$ has strictly improved in $Y$, then condition $1$ is satisfied.
Otherwise, agent $a_{1}$ must have retained the same bundle in $Y$ as in $X$. As no agent envied $a_{1}$ in $X$ who had the same bundle as in $Y$, they continue to not envy $a_{1}$ in $Y$ as $Y\setminus a_{1}\succ X\setminus a_{1}$. Therefore $a_{1}$ is a source in $Y$. Additionally, if agent $a_{1}$ envied some agent in $(\mathbb{B}\cup\mathbb{C})\setminus L$ in $X$, then $a_{1}$ continues to envy the same agent in $Y$ as $Y(\mathcal{N}\setminus L)=X(\mathcal{N}\setminus L)$. ∎
4 agent $a_{1}$ envies someone in $\mathbb{B}\cup\mathbb{C}$
In this section we consider the case when agent $a_{1}$ is a source and envies some agent from $\mathbb{B}\cup\mathbb{C}$.
If $a_{1}$ envies either $b_{1}$ or $c_{1}$, we have the following lemma.
Lemma 7.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If $a_{1}$ is a source in $E_{X}$ and $a_{1}$ envies either $b_{1}$ or $c_{1}$, then there exists an EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
Proof.
Without loss of generality, assume that $a_{1}$ envies $b_{1}$. If $a_{1}$ is the only source in $E_{X}$, we immediately get our desired allocation $Y$ by applying Corollary 2. The remaining case is when there are two sources in $E_{X}$. Thus, $c_{1}$ is also a source in $E_{X}$. Given that some agent must $g$-champion $a_{1}$, we first consider the case where $a_{1}$ self $g$-champions. In this case, we can improve $a_{1}$’s bundle by replacing $X_{a_{1}}$ with the $\mathsf{MES}$ bundle $T_{a_{1}}^{a_{1}}\cup g$. We get $Y$, $\phi(Y)>\phi(X)$.
If, however, $b_{1}$ $g$-champions $a_{1}$, then $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$ is a $\phi$-improving exchange, thus we are done. The only remaining possibility is that $c_{1}$ $g$-champions $a_{1}$. We now consider cases based on who $g$-champions $c_{1}$.
  1. Case 1
    Agent $a_{1}$ $g$-champions $c_{1}$: Then $a_{1}$ and $c_{1}$ form a $g$-champion cycle. Note that none of the leading agents envy $a_{1}$ or $c_{1}$ as they are both sources. In this case, we can apply Lemma 5 to obtain an allocation $Y$ such that $\phi(Y)>\phi(X)$.
  2. Case 2
    Agent $c_{1}$ self $g$-champions: We compute a new Pareto-dominating EFX allocation by replacing $X_{c_{1}}$ with the $\mathsf{MES}$ bundle $T_{c_{1}}^{c_{1}}\cup g$. In this new allocation, $a_{1}$ continues to be a source and still envies $b_{1}$. We start again by applying Lemma 7 to this new allocation $Y$. Since Pareto dominance is a partial order and there are finitely many partial EFX allocations, this procedure must terminate with an allocation $Z$ such that either $Z$ is a complete EFX allocation, or $\phi(Z)>\phi(X)$.
  3. Case 3
    Agent $b_{1}$ $g$-champions $c_{1}$: See figure below for an illustration of this case.
    $a_{1}$ $b_{1}$ $c_{1}$ $g$$g$
    We consider a new allocation by exchanging the top halves as follows:
[…]
    Observe that the bundles of all the three agents $a_{1},b_{1}$, and $c_{1}$ have strictly improved and therefore $\phi(Y)>\phi(X)$. We now check if the new allocation is indeed EFX. If there are no strong envies in $Y$, we are done. So, consider the case when there is at least one strong envy in $Y$. Observer that if there are any strong envies in $Y$, they must be towards either $a_{1}$ or $b_{1}$ or $c_{1}$. We know that no agent strongly envies $a_{1}$ or $b_{1}$ as $a_{1}$ received an old bundle from $X$ and $Y_{b_{1}}$ was an $\mathsf{MES}$ bundle in $X$. Furthermore, agent $b_{1}$ does not envy $c_{1}$. This is because, from Proposition 3 we have $T_{c_{1}}^{b_{1}}>_{b}T_{a_{1}}^{c_{1}}$. Therefore, $X_{c_{1}}=T_{c_{1}}^{b_{1}}\cup B_{c_{1}}^{b_{1}}>_{b}T_{a_{1}}^{c_{1}}\cup B_{c_{1}}^{b_{1}}$. Since $b_{1}$ did not envy $X_{c_{1}}$ in $X$, she does not envy $T_{a_{1}}^{c_{1}}\cup B_{c_{1}}^{b_{1}}$ in $Y$. Thus, the only possible strong envy is from agent $a_{1}$ or other agents in $\mathbb{A}$, towards $c_{1}$. We know that $c_{1}$ does not envy either $a_{1}$ or $b_{1}$. However, $c_{1}$ might envy one of the other agents. We consider the following cases based on possible envies from agent $c_{1}$ in the allocation $Y$.
    1. Case 3.1
      Agent $c_{1}$ envies some $a_{i}\in\mathbb{A}\setminus\{a_{1}\}$: Then, we exchange the bundles of agents $a_{i}$ and $c_{1}$. Note that there may not be envy from $a_{i}$ to $c_{1}$, But since $a_{1}$ envies $c_{1}$, the bundle of $c_{1}$ is better than bundle of $a_{1}$ according to $v_{a}$. Now, all the strong envies are within the groups. Therefore, we apply Lemma 3 to resolve such envies and obtain a new EFX allocation $Z$ such that $\phi(Z)=\phi(Y)>\phi(X)$.
    2. Case 3.2
      Agent $c_{1}$ envies some $b_{j}\in\mathbb{B}\setminus\{b_{1}\}$: In this case, we reallocate as follows: assign $Y_{b_{j}}$ to $c_{1}$, $Y_{c_{1}}$ to $a_{1}$, and $Y_{a_{1}}$ (which is same as $X_{b_{1}}$) back to $b_{1}$. As $a_{1}$ received a bundle which was strongly envied by some agent in $\mathbb{A}$, $a_{1}$’s new bundle is strictly better than $X_{a_{1}}$. Since $c_{1}$ received the bundle that she envied, she also sees a strict improvement.
      However, the value of $b_{1}$’s bundle decreases. Recall that in the previous allocation $X$, we established that $X_{c_{1}}>_{b}T_{a_{1}}^{c_{1}}\cup B_{c_{1}}^{b_{1}}$, therefore $b_{1}$ does not envy $a_{1}$. Furthermore, since $Y_{b_{1}}$ was an $\mathsf{MES}$ bundle relative to $X_{b_{1}}$, agent $b_{1}$ does not strongly envy $c_{1}$. At this point, all remaining strong envies are confined within the groups. Therefore, we apply Lemma 4 to resolve these envies, yielding a new EFX allocation $Z$ such that $\phi(Z)=\phi(Y)>\phi(X)$.
    3. Case 3.3
      Agent $c_{1}$ does not envy anyone from $\mathbb{A}\cup\mathbb{B}$: Let $S\subseteq Y_{c_{1}}$ denote the smallest subset of $Y_{c_{1}}$ such that, according to $c_{1}$’s valuation, $S$ is preferred over the bundles of all agents in $\mathbb{A}\cup\mathbb{B}$. If $a_{1}$ does not strongly envy $S$, then we allocate $S$ to $c_{1}$ and resolve any internal strong envies. This gives us the desired EFX allocation which $\phi$-dominates the previous one.
      Suppose $a_{1}$ strongly envies $c_{1}$ even with the bundle $S$. Then, let $S^{\prime}\subset S$ be a smallest cardinality subset of $S$ that $a_{1}$ weakly envies. Let $T=\max_{c}\{X_{p}\mid p\in\mathbb{A}\cup\mathbb{B}\}$ be the best bundle among the agents in $\mathbb{A}\cup\mathbb{B}$ according to agent $c_{1}$. $T$ can either be a bundle from one of the agents in $\mathbb{A}$ or from $\mathbb{B}$.
      1. Case 3.3.1
        $\mathbf{T=Y_{a_{i}}}$ for some $\mathbf{a_{i}\in\mathbb{A}}$: Then allocate $T$ to $c_{1}$ and $S^{\prime}$ to $a_{i}$. We then resolve internal strong envies within the groups using Lemma 3, yielding a new EFX allocation $Z$ such that $\phi(Z)=\phi(Y)>\phi(X)$.
      2. Case 3.3.2
        $\mathbf{T=Y_{b_{j}}}$ for some $\mathbf{b_{j}\in\mathbb{B}}$ In this case, we ask agent $b_{j}$ to select her most preferred bundle from the agents in $\mathbb{A}$, say from agent $a_{i}$. We then allocate $Y_{a_{i}}$ to $b_{j}$, $Y_{b_{j}}$ to $c_{1}$, and $S^{\prime}$ to agent $a_{i}$. We know that agent $b_{1}$ does not strongly envy any agent from $\mathbb{A}\cup\mathbb{C}$. Thus, we swap the bundles of $b_{j}$ and $b_{1}$, so that now agent $b_{j}$ does not strongly envy anyone from $\mathbb{A}\cup\mathbb{C}$. Consequently, any strong envy towards agents in $\mathbb{A}\cup\mathbb{C}$, if it exists, must now come from agent $b_{1}$.
        In this allocation, agent $c_{1}$ does not envy any agent from $\mathbb{A}\cup\mathbb{B}$ because she chose her most preferred bundle. Agent $a_{1}$’s prior strong envy towards $c_{1}$ is now resolved. Additionally, agent $a_{1}$ does not strongly envy $b_{1}$, as she did not strongly envy $a_{i}$ earlier. Therefore, $a_{1}$ does not strongly envy any agent from $\mathbb{B}\cup\mathbb{C}$.
        Finally, agent $b_{1}$ does not envy any agent in $\mathbb{A}\setminus\{a_{i}\}$, as $b_{j}$ (who has the same valuation as $b_{1}$) selected the best bundle from $\mathbb{A}$. The bundle $Y_{a_{i}}$, which $b_{j}$ picked, is at least as valuable as $Y_{a_{1}}=X_{b_{1}}$. Moreover, recall that with bundle $X_{b_{1}}$, agent $b_{1}$ did not strongly envy $b_{j}$ or any agent in $\mathbb{C}$ in allocation $X$. Therefore, in the new allocation, with a bundle at least as valuable as $X_{b_{1}}$, agent $b_{1}$ does not strongly envy any agent in $\mathbb{C}$. Additionally, we know that $X_{c_{1}}>_{b}T_{a_{1}}^{c_{1}}\cup B_{c_{1}}^{b_{1}}>_{b}S^{\prime}$, so agent $b_{1}$ does not strongly envy $a_{i}$.
        Based on the above observations, agent $b_{1}$ does not strongly envy any agent from $\mathbb{B}\cup\mathbb{A}$. Therefore, we apply Corollary 3 to resolve any internal strong envies, resulting in a new EFX allocation $Z$ such that $\phi(Z)=\phi(Y)>\phi(X)$.∎
Lemma 8.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If agent $a_{1}$ is a source in $E_{X}$ and if $a_{1}$ envies some agent from $\mathbb{B}\cup\mathbb{C}$, then there exists an EFX allocation $Y$ such $\phi(Y)>\phi(X)$.
Proof.
Agent $a_{1}$ envies some agent in $\mathbb{B}\cup\mathbb{C}$. If $a_{1}$ envies either $b_{1}$ or $c_{1}$, we can apply Lemma 7 to get an allocation $Y$ such that $\phi(Y)>\phi(X)$. Thus, for the remainder of this proof, we consider the case where agent $a_{1}$ envies some agent in $(\mathbb{B}\cup\mathbb{C})\setminus\{b_{1},c_{1}\}$.
We now consider various cases based on the number of sources in the envy graph $E_{X}$.
  1. Case 1
    $E_{X}$ has three sources: From Proposition 4, we know that none other than leading agents $L=\{a_{1},b_{1},c_{1}\}$ are the sources in $E_{X}$. We apply Corollary 4 to get an allocation $Y$. If $\phi(Y)>\phi(X)$, we are done. Otherwise, because of Condition $2$ of Corollary 4, we know that $Y\succ X$ and $Y$ satisfies the conditions required to apply Lemma 8. So, we apply Lemma 8 on $Y$. Since Pareto dominance is a partial order and there are finitely many partial EFX allocations, this procedure must terminate with an allocation $Z$ such that either $Z$ is a complete EFX allocation or $\phi(Z)>\phi(X)$.
  2. Case 2
    Agent $a_{1}$ is a single source in $E_{X}$: Then by applying Lemma 2, we get an allocation $Y$ such that $\phi(Y)>\phi(X)$.
  3. Case 3
    $E_{X}$ has two sources: Without loss of generality, assume that $a_{1}$ envies some agent $b_{i}\in\mathbb{B}\setminus\{b_{1}\}$. Furthermore, w.l.o.g assume that there are no self $g$-champions. We know that some agent must $g$-champion $a_{1}$.
    1. Case 3.1
      Agent $b_{1}$ $g$-champions $a_{1}$: Then, $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ forms a $\phi$-improving cycles. Therefore, using Lemma 3, we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
    2. Case 3.2
      Agent $c_{1}$ $g$-champions $a_{1}$: Recall that both $b_{1}$ and $c_{1}$ cannot be sources in $E_{x}$ because of Case $3$ assumption. Furthermore, we have assumed that $a_{1}$ envies neither $b_{1}$ nor $c_{1}$. Therefore, either $b_{1}$ envies $c_{1}$ or $c_{1}$ envies $b_{1}$.
      1. Case 3.2.1
        Agent $b_{1}$ envies $c_{1}$: Here $\langle(a_{1},b_{i}),(b_{1},c_{1}),(c_{1},a_{1})\rangle$ is a $\phi$-improving cycle. Therefore, using Lemma 3, we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
      2. Case 3.2.2
        Agent $c_{1}$ envies $b_{1}$: In this case, we know that someone must $g$-champion agent $c_{1}$.
        1. Case 3.2.2.1
          $b_{1}$ champions $c_{1}$: we construct a new EFX allocation $Y$ by giving the bundle $X_{b_{1}}$ to $c_{1}$ and $T_{c_{1}}^{b_{1}}\cup g$ to agent $b_{1}$. Note that both agent $b_{1}$ and $c_{1}$ have strictly improved. Furthermore, as $T_{c_{1}}^{b_{1}}\cup g$ is a minimally envied subset, no agent strongly envies $b_{1}$. As $X_{b_{1}}$ was not strongly envied in $X$, it cannot be strongly envied in $Y$. Therefore, $Y$ is an EFX allocation such that $Y\succ X$. Also note that agent $a_{1}$ continues to envy agent $b_{i}$. Thus, allocation $Y$ satisfies the conditions required to apply Lemma 8 and $Y\succ X$. So, we apply Lemma 8 on $Y$. Since Pareto dominance is a partial order and there are finitely many partial EFX allocations, this iterative procedure must terminate with an allocation $Z$ such that either $Z$ is a complete EFX allocation or $\phi(Z)>\phi(X)$.
        2. Case 3.2.2.2
          Agent $a_{1}$ $g$-champions $c_{1}$: Then, we exchange the top halves of $X_{a_{1}}$ and $X_{c_{1}}$ to get a new allocation $Y$, where the bundles of $a_{1},b_{1}$ and $c_{1}$ are as follows:
[…]
          Applying Proposition 3, we know that $T_{c_{1}}^{a_{1}}>_{a}T_{a_{1}}^{c_{1}}$. Therefore, $Y_{a_{1}}>_{a}X_{a_{1}}>_{a}Y_{c_{1}}$. That is, agent $a_{1}$ has strictly improved and no agent from $\mathbb{A}$ strongly envies $c_{1}$ in $Y$. Analogously, $Y_{c_{1}}>_{c}X_{c_{1}}>_{c}Y_{a_{1}}$. That is, agent $c_{1}$ has strictly improved and no agent from $\mathbb{C}$ strongly envies $a_{1}$ in $Y$. Since agent $b_{1}$ was envying neither $a_{1}$ nor $c_{1}$ in $X$, $b_{1}$ can envy at most one of $a_{1},c_{1}$ in $Y$. Without loss of generality, assume that $b_{1}$ strongly envies $a_{1}$. We then replace the bottom half $B_{a_{1}}^{c_{1}}$ of $a_{1}$ with $g$ to get $Z$. The bundles of $a_{1},b_{1},$ and $c_{1}$ in $Z$ are as follows:
[…]
          As $T_{c_{1}}^{a_{1}}\cup g$ is a minimally envied subset, no one strongly envies agent $a_{1}$. Furthermore, $Z_{a_{1}}>_{a}X_{a_{1}}>_{a}$, therefore agent $a_{1}$ is strictly better off compared to $X$ and $a_{1}$ does not strongly envy $c_{1}$. Therefore, we have an EFX allocation $Z$ such that $\phi(Z)>\phi(X)$. ∎
5 Agent $a_{1}$ envies no one in $\mathbb{B}\cup\mathbb{C}$
Here we consider a partial EFX allocation $X$, where agent $a_{1}$ is a source and $a_{1}$ does not envy anyone in $\mathbb{B}\cup\mathbb{C}$.
Lemma 9.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If $a_{1}$ is a source in $E_{X}$ and $a_{1}$ envies no agent from $\mathbb{B}\cup\mathbb{C}$, then there exists an EFX allocation $Y$ such $\phi(Y)>\phi(X)$.
Proof.
We consider all possible cases based on the number of sources in $E_{X}$.
  1. Case 1.
    If $a_{1}$ is the only source in $E_{X}$, we apply Corollary 2 to get an EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
  2. Case 2.
    Suppose there are three sources in $E_{X}$. That is, $L=\{a_{1},b_{1},c_{1}\}$ is the set of sources in $E_{X}$. We apply Corollary 4 on $X$ to get an allocation $Y$, $Y\succ X$. If $\phi(Y)>\phi(X)$, we are done. Otherwise, we know that $a_{1}$ is a source in $Y$. If $a_{1}$ envies some agent in $\mathbb{B}\cup\mathbb{C}$, we apply Lemma 8 on $Y$ to get an allocation $Z$, such that $\phi(Z)>\phi(Y)=\phi(X)$. If $a_{1}$ envies no agent in $\mathbb{B}\cup\mathbb{C}$, we apply Lemma 9 on $Y$, as $Y$ satisfies all the conditions to apply Lemma 9 and $Y\succ X$. Since Pareto dominance is a partial order and there are finitely many partial EFX allocations, this iterative procedure must terminate with an EFX allocation $Z$ such that either $Z$ is a complete allocation or $\phi(Z)>\phi(X)$.
  3. Case 3.
    There are two sources in $E_{X}$. We know that $a_{1}$ is a source. Without loss of generality, let $b_{1}$ be the other source. Therefore $c_{1}$ is not a source and hence it must be envied by some leading agent. Since $a_{1}$ does not envy anyone, $b_{1}$ envies $c_{1}$. We now consider the cases depending on which agent $g$-champions $b_{1}$.
    1. Case 3.1.
      Suppose agent $b_{1}$ self $g$-champions (see the figure below). We then construct a new Pareto dominating EFX allocation $Y$ by replacing $X_{b_{1}}$ with $\mathsf{MES}_{b}(X_{b_{1}}\cup g)$. Since this is a strict improvement for agent $b_{1}$, she continues to not envy $a_{1}$ even in $Y$. No one else envies $a_{1}$ as there was no change in their bundles. Thus $a_{1}$ is a source in $Y$. If $a_{1}$ envies $b_{1}$, we apply Lemma 8 to get a $\phi$ improvement. Otherwise, we apply Lemma 9 again on $Y$. As there are finitely many EFX allocations, this Pareto improvement cannot go on for ever.
      $a_{1}$ $b_{1}$ $c_{1}$ $g$
    2. Case 3.2.
      If $c_{1}$ $g$-champions $b_{1}$ (See the figure below), then $(b_{1},c_{1},b_{1})$ is a Pareto improving cycle. Eliminating this cycle gives an allocation $Y$ such that $Y\succ X$ and $a_{1}$ is a source in $E_{Y}$ and $a_{1}$ does not envy any agent in $\mathbb{B}\cup\mathbb{C}$. We apply Lemma 9 again on $Y$. As there are finitely many EFX allocations, this Pareto improvement cannot go on for ever.
      $a_{1}$ $b_{1}$ $c_{1}$ $g$
    3. Case 3.3.
      If $a_{1}$ $g$-champions $b_{1}$, then we give an algorithm (Algorithm 1) to get an allocation $Y$ such that $\phi(Y)>\phi(X)$. Lemma 13 proves the correctness of Algorithm 1. ∎
5.1 Agent $a_{1}$ is a source, $a_{1}$ envies no one in $\mathbb{B}\cup\mathbb{C}$ but $g$-champions $b_{1}$
In this section, we consider the case when we have a partial EFX allocation $X$ with an unallocated item $g$, in which, agent $a_{1}$ is a source, and $a_{1}$ does not envy anyone in $\mathbb{B}\cup\mathbb{C}$ but $a_{1}$ $g$-champions $b_{1}$. See the figure below.
$a_{1}$ $b_{1}$ $c_{1}$ $g$
The pseudocode for our algorithm to handle this case is presented as Algorithm 1. It takes $X$ as input and outputs another EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
5.1.1 Overview of Algorithm 1
In this section, we give an overview of the algorithm. Please refer to Algorithm 1 the pseudo-code.
High-level idea: Recall that $a_{1}$ does not envy $b_{1}$ but $a_{1}$ $g$-champions $b_{1}$ with respect to the EFX allocation $X$. So the bundle $T_{b_{1}}^{a_{1}}\cup g$ is such that $a_{1}$ does not strongly envy anyone after receiving $T_{b_{1}}^{a_{1}}\cup g$ and in fact, this hold for any $a_{i}$, with respect to $X$. Further, $T_{b_{1}}^{a_{1}}\cup g>_{a}X_{a_{1}}$. Also, no one strongly envies $T_{b_{1}}^{a_{1}}\cup g$ with respect to $X$. If $T_{b_{1}}^{a_{1}}\cup g$ can be allotted to $a_{1}$, and a suitable replacement bundle for $b_{1}$ can be found, then this will result in an EFX allocation with improved value of $\phi$. This is roughly the goal of Algorithm 1. Temporarily, we allot $T_{b_{1}}^{a_{1}}\cup g$ to $b_{1}$. Clearly, the strong envy can only be from $b_{1}$ to other agents. This invariant, called Invariant $1$, is maintained throughout Algorithm 1.
Invariant 1.
In the allocation $X$, only agent $b_{1}$ may have strong envy towards others.
Notation: At the beginning of Algorithm 1, define $O_{b}$, the comparison set for $\mathbb{B}$, as the set of all the bundles allocated to agents in $\mathbb{A}\cup\mathbb{C}$, together with $T_{b_{1}}^{a_{1}}\cup g$. During the course of Algorithm 1, some of the bundles in $O_{b}$ will be replaced by one of their proper subsets, called a virtual bundle. At any point in the algorithm, denote $P=\max_{b}O_{b}$. Similarly, define $O_{c}$, the comparison set for $\mathbb{C}$ as the bundles belonging to agents in $\mathbb{A}\cup\mathbb{B}$. Define $Q=\max_{c}(O_{c}\cup\{P\})$.
  1. Case 1
    $\mathbf{P\in\{X_{a_{i}}\mid a_{i}\in\mathbb{A}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}}$: See Figure 2. If $P=X_{i}$ for some $i$, then $P$ can be allocated to $b_{1}$ and $T_{b_{1}}^{a_{1}}\cup g$ to $a_{i}$, followed by internal sorting of bundles of groups $\mathbb{A}$ and $\mathbb{B}$. Clearly, after receiving $P$, $b_{1}$ has no strong envy towards any agent in $\mathbb{A}\cup\mathbb{C}$. Possible strong envy from $b_{1}$ to other agents in $\mathbb{B}$ can be resolved using Lemma 4 to get an EFX allocation. If $P=X_{a_{1}}$ then $a_{1}$ gets a better bundle than $X_{a_{1}}$, namely $T_{b_{1}}^{a_{1}}\cup g$, and hence $\phi$ improves. If $P=X_{a_{i}}$ for $i\neq 1$, then $X_{a_{1}}$ still remains the minimum valued bundle in group $\mathbb{A}$. However, $a_{1}$ envies $P$ since $P=X_{a_{i}}>_{a}X_{a_{1}}$. Thus we can apply Lemma 8 to get an EFX allocation with a higher value of $\phi$. A similar argument applies if $P=T_{b_{1}}^{a_{1}}\cup g$.
[…]
    Figure 2: Green edge points to $P=X_{a_{i}}=\max_{b}O_{b}$
  2. Case 2
    $\mathbf{P\subseteq X_{c_{i}}}$ for some $\mathbf{i}$ and $\mathbf{Q\in\{X_{a_{j}}\mid a_{j}\in\mathbb{A}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}}$: See Figure 3. If $Q=X_{a_{j}}$ for some $j$, a similar improvement in $\phi$ can be obtained by giving $T_{b_{1}}^{a_{1}}\cup g$ to $a_{j}$, $P$ to $b_{1}$, and $Q$ to $c_{i}$. Note that this might need resolution of intra-group strong envies, which can be done by Lemma 4, and also sorting the bundles internally within each group so that the leading agents retain minimum valued bundles in their respective groups. A similar $\phi$ improvement can be made if $Q=T_{b_{1}}^{a_{1}}\cup g$ by using Lemma 8.
[…]
    Figure 3: Green edge points to $P$ and cyan edge to $Q$
  3. Case 3
    Both the above conditions do not hold: It must be the case that either $P\subseteq X_{c_{i}}$ and $Q=X_{b_{j}}$ for some $i,j$, $j\neq 1$ or $Q=P$. However, even in the former case, we cannot simply swap $P$ and $Q$ between $c_{i}$ and $b_{j}$, since it is not clear whether this procedure will then terminate.
    1. Case 3a
      $\mathbf{Q=X_{b_{j}}}$ for some $\mathbf{b_{j}\in\mathbb{B}\setminus\{b_{1}\}}$: In this case, the bundle $X_{c_{i}}$ in $O_{b}$ is replaced by $X_{c_{i}}\setminus h$, where $h$ is the least valued good in $X_{c_{i}}$ according to $v_{b}$. Note that this operation does not alter the actual allocation. We refer to this update of $O_{b}$ as a virtual update. The algorithm starts a new iteration with this updated set $O_{b}$. The idea is that the virtual update can happen at most once for each bundle of group $\mathbb{C}$. At some point, either $b_{1}$ chooses such a virtual bundle as $\max_{b}O_{b}$, or chooses some bundle of group $\mathbb{A}$, satisfying Case 1 above.
    2. Case 3b
      $\mathbf{Q=P=X_{c_{i}}}$: Then, agents $b_{1}$ and $c_{i}$ compete for $P$ with $O_{b}$ and $O_{c}$ as their respective comparison sets. Each of $b_{1}$ and $c_{i}$ picks the minimum cardinality subsets, say $P_{b}$ and $P_{c}$ of $P$ respectively, such that, given this subset, they will have no envy towards any bundle in their respective comparison sets. We call $c_{i}$ to be the winner if $|P_{c}|\leq|P_{b}|$, and we call $P_{c}$ as the winning subset.
      1. Case 3b(i)
        $\mathbf{c_{i}}$ is a winner and $\mathbf{P_{c}\subsetneq X_{c_{i}}}$: We then update the allocation $X$ by allocating $P_{c}$ to $c_{i}$, and adding the goods in $X_{c_{i}}\setminus P_{c}$ to the set $U$ of unallocated goods. By the definition of a winning subset, when $c_{i}$ is given $P_{c}$, agent $c_{i}$ can have possible strong envies towards only the agents in $\mathbb{C}$. Such strong envies are resolved using Lemma 4. At this point, we have a new allocation that satisfies Invariant 1 with a strictly larger number of unallocated goods. Therefore, the comparison sets $O_{b}$ and $O_{c}$ are updated to reflect the new allocation: $O_{b}\leftarrow\{X_{\alpha}\mid\alpha\in\mathbb{A}\cup\mathbb{C}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$, and $O_{c}\leftarrow\{X_{\alpha}\mid\alpha\in(\mathbb{A}\cup\mathbb{B})\}$. The algorithm then proceeds with a new iteration.
      2. Case 3b(ii)
        $\mathbf{c_{i}}$ is a loser, or $\mathbf{P_{c}=P_{b}=X_{c_{i}}}$. In this case, a virtual update is done. That is, the bundle $X_{c_{i}}$ in $O_{b}$ is replaced by $X_{c_{i}}\setminus h$, where $h$ is the least valued good in $X_{c_{i}}$ according to $v_{b}$. This does not alter the actual allocation $X$. The algorithm continues with the modified set $O_{b}$. As in Case 3Case 3a above, this cannot go on indefinitely.
    3. Case 3c
      $\mathbf{P}$ is a virtual bundle i.e. $\mathbf{P\subsetneq X_{c_{i}}}$ for some $\mathbf{c_{i}\in\mathbb{C}}$: This is considered on Line 1 of Algorithm 1. This can happen since $O_{b}$ might have some virtual bundles from earlier iterations. Thus $P=X_{c_{i}}\setminus h$. Then either $Q=P$ or $Q$ is a bundle belonging to group $\mathbb{A}$ or $\mathbb{B}$. If $Q=P$, we simply allot $Q$ to $c_{i}$ and add $h=X_{c_{i}}\setminus Q$ to the set of unallocated goods $U$. If $Q=X_{b_{j}}$ for some $j$, we allot $X_{c_{i}}\setminus h$ to $b_{j}$, $X_{b_{j}}$ to $c_{i}$, and add $h$ to $U$. In both these cases, $h$ gets included in the pool of unallocated items and the pool of unallocated items grows. We update $O_{b}$ and $O_{c}$ and start a new iteration. Also, it is ensured that if an agent in $\mathbb{B}$ does not envy a virtual bundle, then she will not strongly envy the corresponding real bundle, thus $b_{1}$ has no strong envy towards agents in $\mathbb{A}\cup\mathbb{C}$. We will prove it in Lemma 11 later.
The correctness of Case 1 and Case 2 above is proved respectively in Part 1 and Part 2 of Lemma 10 below.
Input: An EFX allocation $X$ such that $a_{1}$ $g$-champions $b_{1}$ and $a_{1}$ does not envy $b_{1}$
Output: An EFX allocation $Y$ such that $\phi(Y)>\phi(X)$
1
2Allocate $\{T_{b_{1}}^{a_{1}}\cup g\}$ to agent $b_{1}$
3 Let $O_{b}\leftarrow\{X_{\alpha}\mid\alpha\in\mathbb{A}\cup\mathbb{C}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$ and let $O_{c}\leftarrow\{X_{\alpha}\mid\alpha\in(\mathbb{A}\cup\mathbb{B})\}$
4 Let $U$ be the set of unallocated goods in $X$
5
6while true do
7       Let $P\leftarrow\max_{b}O_{b}$
8       if $P\in\{X_{a_{i}}\mid a_{i}\in\mathbb{A}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$ then
9             Apply Lemma 10 (Part $(i)$), obtaining an EFX allocation $Y$, such that $\phi(Y)>\phi(X)$
10             return $Y$
11
12
13      else  // $P\subseteq X_{c_{i}}$ for some $c_{i}\in\mathbb{C}$
14
15            Let $Q\leftarrow\max_{c}(O_{c}\cup P)$
16             if $Q\in\{X_{a_{i}}\mid a_{i}\in\mathbb{A}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$ then
17                   Apply Lemma 10 (Part $(ii)$), obtaining an EFX allocation $Y$, such that $\phi(Y)>\phi(X)$
18                   return $Y$
19
20
21            else if $P=X_{c_{i}}$ for some $c_{i}\in\mathbb{C}$ then
22                   if $Q=P$ then
23                         $b_{1}$ and $c_{i}$ compete for $P$
24                        if $c_{i}$ is a winner and the winning subset $P_{c}\subsetneq P$ then
25                               In $X$, replace $X_{c_{i}}$ with the winning subset $P_{c}\subsetneq X_{c_{i}}$
26                               $U\leftarrow U\cup(X_{c_{i}}\setminus P)$
27                               Resolve internal strong envies in $\mathbb{C}$ (Lemma 4)
28                               Update: $O_{c}\leftarrow\{X_{\alpha}\mid\alpha\in(\mathbb{A}\cup\mathbb{B})\}$ and $O_{b}\leftarrow\{X_{\alpha}\mid\alpha\in\mathbb{A}\cup\mathbb{C}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$
29
30                        else
31                               In $O_{b}$, replace $X_{c_{i}}$ with $X_{c_{i}}\setminus h$, where $h=\min_{b}X_{c_{i}}$
32
33
34                  else if $Q\in\{X_{b_{j}}\mid b_{j}\in\mathbb{B}\setminus\{b_{1}\}\}$ then
35                         In $O_{b}$, replace $X_{c_{i}}$ with $X_{c_{i}}\setminus h$, where $h=\min_{b}X_{c_{i}}$
36
37
38            else if $P\subsetneq X_{c_{i}}$ for some $c_{i}\in\mathbb{C}$ then
39                   if $Q=P$ then
40                         In $X$, replace $X_{c_{i}}$ with $P$
41
42
43                  else  // $Q=X_{b_{j}}$ for some $b_{j}\in\mathbb{B}\setminus\{b_{1}\}$
44
45                        Modify $X$: allocate $P$ to $b_{j}$, $X_{b_{j}}$ to $c_{i}$
46                         Resolve $b_{j}$ to $b_{k}\in\mathbb{B}\setminus b_{1}$ strong envies by replacing $X_{b_{k}}$ with $\mathsf{MES}_{b}(X_{b_{k}},P)$
47                        Resolve internal strong envies in $\mathbb{C}$ (Lemma 4)
48                         $U\leftarrow U\cup(X_{c_{i}}\setminus P)$
49                         Update: $O_{b}\leftarrow\{X_{\alpha}\mid\alpha\in\mathbb{A}\cup\mathbb{C}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$ and $O_{c}\leftarrow\{X_{\alpha}\mid\alpha\in(\mathbb{A}\cup\mathbb{B})\}$
50
51
52
53
Algorithm 1 Multi-Round-Competition
Lemma 10.
Let $X$ be a partial allocation with an unallocated good $g$, where all strong envies in $X$, if any, originate from agent $b_{1}$. Suppose agent $a_{1}$ is a source in $E_{X}$, does not envy $b_{1}$, but $g$-champions $b_{1}$, partitioning $X_{b_{1}}$ as $X_{b_{1}}=T_{b_{1}}^{a_{1}}\uplus B_{b_{1}}^{a_{1}}$. Define:
$O_{b}$ =$=\{X_{\alpha}\mid\alpha\in\mathbb{A}\cup\mathbb{C}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$
$O_{c}$ =$=\{X_{\alpha}\mid\alpha\in(\mathbb{A}\cup\mathbb{B})\setminus\{b_{1}\}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$
If either of the following conditions is satisfied, then there exists an EFX allocation $Y$ (possibly partial) such that $\phi(Y)>\phi(X)$.
  1. 1.
    There exists a bundle $P\in\{X_{\alpha}\mid\alpha\in\mathbb{A}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$ such that, given $P$, agent $b_{1}$ does not strongly envy any bundle in $O_{b}$,
  2. 2.
    There exists a bundle $P\subseteq X_{c_{i}}$ for some $c_{i}\in\mathbb{C}$, such that given $P$, agent $b_{1}$ does not strongly envy any bundle in $O_{b}$, and there exists some bundle $Q\in\{X_{\alpha}\mid\alpha\in\mathbb{A}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}$ such that with $Q$ agent $c_{i}$ does not strongly envy any bundle in $O_{c}\cup\{P\}$.
Proof.
  1. (i)
    If $P=\{T_{b_{1}}^{a_{1}}\cup g\}$, then we allocate $\{T_{b_{1}}^{a_{1}}\cup g\}$ to $b_{1}$. We know that with $P$, agent $b_{1}$ does not strongly envy anyone from $\mathbb{A}\cup\mathbb{C}$. Resolve internal envies in $\mathbb{B}$, if any, using Lemma 4. We now have an EFX allocation $Y$ such that agent $a_{1}$ envies $b_{1}$. Therefore, by applying Lemma 8 on $Y$, we get an EFX allocation $Z$ such that $\phi(Z)>\phi(X)$.
    If $P=X_{a_{1}}$, we then give $P$ to agent $b_{1}$ and $\{T_{b_{1}}^{a_{1}}\cup g\}$ to agent $a_{1}$. Resolve internal envies in $\mathbb{B}$, if any, using Lemma 4. We now have an EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
    Finally, if $P=X_{a_{i}}$ for some $a_{i}\in\mathbb{A}\setminus a_{1}$, then give $P$ to agent $b_{1}$ and $\{T_{b_{1}}^{a_{1}}\cup g\}$ to agent $a_{i}$. Resolve internal envies in $\mathbb{B}$ or $\mathbb{A}$, if any, using Lemma 4. We now have a EFX allocation $Y$ such that agent $a_{1}$ envies $b_{1}$. Therefore, by applying Lemma 8 on $Y$, we get an EFX allocation $Z$ such that $\phi(Z)>\phi(X)$.
  2. (ii)
    Let $P\subseteq X_{c_{i}}$ for some agent $c_{i}\in\mathbb{C}$. If $Q=\{T_{b_{1}}^{a_{1}}\cup g\}$, we allocate $P$ to $b_{1}$ and $Q$ to agent $c_{i}$. Else if $Q=X_{a_{j}}$ for some $a_{j}\in\mathbb{A}\setminus a_{1}$ we allocate $P$ to $b_{1}$, $Q$ to $c_{i}$ and $\{T_{b_{1}}^{a_{1}}\cup g\}$ to agent $a_{j}$. In either case, we resolve internal envies in within the groups, if any, using Lemma 4. We now have a EFX allocation $Y$ such that agent $a_{1}$ envies $c_{i}$. Therefore, by applying Lemma 8 on $Y$, we get an EFX allocation $Z$ such that $\phi(Z)>\phi(X)$.
    If $Q=X_{a_{1}}$, then we allocate $P$ to $b_{1}$,$Q$ to $c_{i}$, and $\{T_{b_{1}}^{a_{1}}\cup g\}$ to agent $a_{1}$. Resolve internal envies in within the groups, if any, using Lemma 4. We get an EFX allocation $Y$ such that $\phi(Y)>\phi(X)$. ∎
Lemma 11.
In Algorithm 1, if $b_{1}$ is allocated a virtual bundle from $O_{b}$, then $b_{1}$ does not strongly envy anyone from $\mathbb{A}\cup\mathbb{C}$. In fact, the same holds for any $b_{j}\in\mathbb{B}$.
Proof.
A virtual bundle appears in $O_{b}$ when Case 3Case 3a holds i.e. $\max_{b}O_{b}=X_{c_{i}}$ for some $i$ and $\max_{c}O_{c}=X_{b_{j}}$ for some $j\neq 1$, or Case 3(Case 3b)Case 3b(ii) holds, when $P=Q=X_{c_{i}}$ and either $b_{1}$ has a smaller winning subset than $c_{i}$ i.e. $|P_{b}|<|P_{c}|$ or both have $X_{c_{i}}$ as the winning subset i.e. $P_{b}=P_{c}=X_{c_{i}}$. When $P=\max_{b}O_{b}$ is a virtual bundle, then for any non-virtual i.e. real bundle $S\in O_{b}$, $P>_{b}S$ and hence $b_{1}$ does not envy the agent holding $S$. For any virtual bundle $S=X_{c_{i}}\setminus h\in O_{b}$, $S\neq P$, $P>_{b}X_{c_{i}}\setminus h$ for minimum valued item $h\in X_{c_{i}}$ according to $v_{b}$. Hence $b_{1}$ might envy the real bundle $X_{c_{i}}$ but it is not a strong envy. Since $b_{j}$ and $b_{1}$ have the same valuation $v_{b}$, the lemma holds for any $b_{j}\in B$.∎
To prove the correctness of Case 3, one needs to show that Invariant 1 is maintained as long as Algorithm 1 remains in Case 3, and that the algorithm exits from Case 3 in a finite number of iterations.
Lemma 12.
At any point during the course of Algorithm 1, if Case 3 holds, then in a finite number of iterations, the algorithm exits from Case 3, and further, Invariant 1 is maintained as long as the algorithm remains in Case 3.
Proof.
In each iteration of Algorithm 1, whenever Case 3 holds, either the pool of unallocated items increases or the number of virtual bundles in $O_{b}$ grows. The former happens in cases Case 3(Case 3b)Case 3b(i) and Case 3Case 3c whereas the later happens in cases Case 3Case 3a and Case 3(Case 3b)Case 3b(ii). Since neither of these can continue beyond a finite number of steps (at most as many steps as the number of items), Algorithm 1 exits Case 3 in a finite number of steps.
It can be seen that Invariant 1 holds as long as Algorithm 1 continues in Case 3. When the allocation is modified in Case 3 (Line 1 of Algorithm 1), $b_{j},j\neq 1$ gets a virtual bundle $P=X_{c_{i}}\setminus h$. By Lemma 11, on allocating $P$, $b_{j}$ does not strongly envy anyone from $\mathbb{A}\cup\mathbb{C}$. Similarly, by the choice of $Q$, $c_{i}$ does not strongly envy anyone from $\mathbb{A}\cup\mathbb{C}$. No agent except possibly $b_{1}$ strongly envies $b_{j}$ since they did not strongly envy $c_{i}$ when $c_{i}$ had $X_{c_{i}}$. The potential strong envies from $b_{j}$ towards other agents in $\mathbb{B}$ and from $c_{i}$ to other agents in $\mathbb{C}$ are resolved using Lemma 4. Thus only $b_{1}$ might strongly envy, and hence Invariant 1 continues to hold.∎
The following lemma completes the correctness proof of Algorithm 1.
Lemma 13.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. Let agent $a_{1}$ be a source in $E_{X}$ and does not envy agent $b_{1}$, but $g$-champions $b_{1}$. Then, with $X$ as input, Algorithm 1 terminates with an EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
Proof.
If Case 1 holds at any point, the algorithm, on Line 1 returns an EFX allocation $Y$ with $\phi(Y)>\phi(X)$ by Part 1 of Lemma 10. If Case 2 holds, then on Line 1, an EFX allocation $Y$ with $\phi(Y)>\phi(X)$ is returned by Part 2 of Lemma 10. So the algorithm successfully terminates whenever one of these two cases hold.
Note that the algorithm terminates if and only if it reaches Line 1 or Line 1 i.e. if and only if Case 1 or Case 2 hold. So the while loop on Line 1 can go on as long as the algorithm continues in Case 3. As proved in Lemma 12, Case 3 can hold for only a finite number of iterations. So the algorithm must enter Case 1 or Case 2 in a finite number of iterations, and must terminate with an EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.∎
From Lemma 8 and Lemma 9, we derive the following corollary.
Corollary 5.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If $a_{1}$ is a source in $E_{X}$, then there exists an EFX allocation $Y$ such $\phi(Y)>\phi(X)$.
6 Agent $a_{1}$ is Not a Source in $E_{X}$
When $a_{1}$ is not a source in $E_{X}$, someone must envy $a_{1}$. By Proposition 4, one of $b_{1}$ or $c_{1}$ must envy $a_{1}$. In addition to this, $a_{1}$ may or may not envy anyone from $\mathbb{B}\cup\mathbb{C}$. In Lemma 14, we consider the case when $a_{1}$ envies someone from $\mathbb{B}\cup\mathbb{C}$. The other case when $a_{1}$ does not envy anyone from $\mathbb{B}\cup\mathbb{C}$ is considered in Lemma 15. In each of these cases, we give a construction of another EFX allocation that has a higher value of the potential function.
Lemma 14.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If $a_{1}$ is a not a source in $E_{X}$, and if $a_{1}$ envies some agent in $\mathbb{B}\cup\mathbb{C}$, then there exists an EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
Proof.
since $a_{1}$ is not a source, either $b_{1}$ or $c_{1}$ must envy $a_{1}$ by Proposition 4. Without loss of generality, assume $b_{1}$ envies $a_{1}$.
  1. 1
    Agent $a_{1}$ envies some agent $b_{j}\in\mathbb{B}$: Then, $\langle(a_{1},b_{j}),(b_{1},a_{1})\rangle$ is a $\phi$-improving exchange. Therefore, using Lemma 3 we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
  2. 2
    Agent $a_{1}$ envies some $c_{j}\in\mathbb{C}$: If $c_{1}$ envies $a_{1}$, then $\langle(a_{1},c_{j}),(c_{1},a_{1})\rangle$ is a $\phi$-improving exchange. Thus we are done. Consider the case when $c_{1}$ does not envy $a_{1}$. We know that some leading agent must $g$-champion $b_{1}$. We consider all possible cases based on who $g$-champions $b_{1}$.
    1. 2.1
      Agent $b_{1}$ self $g$-champions: We make a Pareto improvement by replacing $X_{b_{1}}$ with $\mathsf{MES}_{b}(X_{b_{1}}\cup g,X_{b_{1}})$. Following this, the bundles in $\mathbb{B}$ are re-sorted so that $b_{1}$ holds the least valuable bundle. Call this allocation $Y$. If in $Y$, $b_{1}$ does not envy $a_{1}$, then $a_{1}$ is a source in $E_{Y}$. Therefore, applying Corollary 5, we get a new EFX allocation $Z$ such that $\phi(Z)>\phi(Y)=\phi(X)$. Otherwise, if $b_{1}$ envies $a_{1}$ in $Y$, then we apply Lemma 14 again on $Y$. As there are finitely many EFX allocations, this Pareto improvement cannot go on forever.
    2. 2.2
      Agent $a_{1}$ $g$-champions $b_{1}$: Then $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$ is a $\phi$-improving exchange. Therefore, applying Lemma 3, we get new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
    3. 2.3
      Agent $c_{1}$ $g$-champions $b_{1}$: Then we have $\langle(a_{1},c_{j}),(c_{1},b_{1}),(b_{1},a_{1})\rangle$ as a $\phi$-improving exchange. Thus, we get new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.∎
Lemma 15.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If $a_{1}$ is a not a source in $E_{X}$, and if $a_{1}$ envies no agent in $\mathbb{B}\cup\mathbb{C}$, then there exists an EFX allocation $Y$ such that either $Y\succ X$ or $\phi(Y)>\phi(X)$.
Proof.
We know that either $b_{1}$ or $c_{1}$ must envy $a_{1}$. Without loss of generality, assume $b_{1}$ envies $a_{1}$. Some leading agent must $g$-champion $b_{1}$. If $b_{1}$ self $g$-champions, then we can make Pareto improvement, thus we are done. If $a_{1}$ $g$-champions $b_{1}$, we have a $\phi$-improving exchange $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$. Therefore, consider the case when agent $c_{1}$ $g$-champions $b_{1}$.
We know that some leading agent must $g$-champion $c_{1}$. If $c_{1}$ self $g$-champions, then we can make Pareto improvement. Suppose $b_{1}$ $g$-champion $c_{1}$. If either $c_{1}$ envies $b_{1}$ or $b_{1}$ envies $c_{1}$, then $(c_{1},b_{1},c_{1})$ is a Pareto improving cycle, thus we make a Pareto improvement. If not, we have $b_{1}$ and $c_{1}$ $g$-championing each other and $a_{1}$ envies neither $b_{1}$ nor $c_{1}$. We apply Lemma 5 to get Pareto improvement.
The only case remaining is when agent $a_{1}$ $g$-champions $c_{1}$. In this case, we ask $c_{1}$ to pick the best bundle from agents in $\mathbb{A}\cup\mathbb{B}$.
  1. Case 1
    Agent $c_{1}$ picks a bundle $X_{b_{j}}$ from an agent $b_{j}\in\mathbb{B}$: Then, we construct a new allocation $Y$, by allocating $X_{b_{j}}$ to $c_{1}$, $T_{c_{1}}^{a_{1}}\cup g$ to agent $a_{1}$, and $X_{a_{1}}$ to $b_{j}$. Here, agent $a_{1}$ strictly improves. In this new allocation, no agent strongly envies any of $a_{1},b_{j}$ and $c_{1}$ as the bundles that these agents received are either $\mathsf{MES}$ bundles or complete bundles in the old allocation $X$. Agent $b_{j}$ does not strongly envy any agent as $v_{b}(Y_{b_{j}})\geq v_{b}(Y_{b_{1}})$ and agent $b_{1}$ does not strongly envy any agent. As $a_{1}$ received a strict improvement, she does not strongly envy any agent. Finally, we know that $c_{1}$ does not envy any agent from $\mathbb{B}\cup\mathbb{A}$ as it chose the best bundle. If $c_{1}$ strongly envies any agent in $\mathbb{C}$, we resolve them using Lemma 4. We now have $\phi(Y)>\phi(X)$.
  2. Case 2
    Agent $c_{1}$ picks a bundle $X_{a_{1}}$: Then, to construct a new allocation $Y$, allocate $X_{a_{1}}$ to $c_{1}$ and $T_{c_{1}}^{a_{1}}\cup g$ to $a_{1}$. Again, note that if there are any strong envies in $Y$, they must be from agent $c_{1}$ to other agents in $\mathbb{C}$. We resolve such strong envies using Lemma 4. We have $\phi(Y)>\phi(X)$.
  3. Case 3
    Agent $c_{1}$ picks $X_{a_{j}}$, $a_{j}\in\mathbb{A}\setminus\{a_{1}\}$: Then, to construct a new allocation $Y$, allocate $X_{a_{j}}$ to $c_{1}$ and $T_{c_{1}}^{a_{1}}\cup g$ to $a_{j}$. Note that agent $a_{j}$ does not strongly envy anyone as $v_{a}(Y_{a_{j}})>v_{a}(Y_{a_{1}})=v_{a}(X_{a_{1}})$ and agent $a_{1}$ does not strongly anyone. If there are any strong envies in $Y$, they must be from agent $c_{1}$ to other agents in $\mathbb{C}$. We resolve such strong envies using Lemma 4. In $Y$, agent $b_{1}$ continues to envy $a_{1}$ as their bundles did not change. Agent $a_{1}$ envies $c_{1}$ as $a_{1}$ previously envied $a_{j}$. Therefore, we apply Lemma 14 on $Y$ to get a new allocation $Z$ such that $\phi(Z)>\phi(Y)=\phi(X)$.∎
From Lemma 14 and Lemma 15, we get the following corollary.
Corollary 6.
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If $a_{1}$ is a not a source in $E_{X}$, then there exists an EFX allocation $Y$ such either $Y\succ X$ or $\phi(Y)>\phi(X)$.
Therefore, Lemma 8 and Lemma 9 completes the proof of Theorem 1. Hence, EFX exists for three types of agents.
7 Conclusion
In this paper, we proved the existence of EFX allocations for instances with three types of agents, where agents of the same type have identical additive valuations. This result extends the prior work on EFX allocations for three agents [efx_3] and for two types of agents [mahara].
The techniques developed in this work may provide insights into more general settings. A natural direction for future research is to extend these results to agents with more general valuation functions, beyond the additive case. Another long standing open question is whether EFX allocations can be guaranteed for instances involving four or more agents.
Given the progress made here, we propose the following conjecture as a potential pathway for further exploration:
Conjecture 1.
If EFX allocations exist for $k$ agents with additive valuations, for some constant $k$, then EFX allocations exist for any number of agents with at most $k$ distinct types.
This conjecture, if proven, would broaden the scope of EFX allocations, particularly in practical settings where large groups of agents share similar preferences.